### v2

# Offline Dedicated Quantum Attacks on Block Ciphers Based on Two Parallel Permutation-Based Pseudorandom Functions

Abstract
Quantum cryptanalysis is essential for evaluating the security of cryptographic systems against the threat of quantum computing. Recently, Shi et al. introduced the dedicated quantum attack on XOR-type function that greatly reduces the required resources (including circuit depth, width, and the number of gates) compared to the parallel Grover-meets-Simon algorithm. Here, our contribution is in two aspects. On the one hand, we discover new cryptographic structures amenable to this attack: PolyMAC and block ciphers based on two parallel permutation-based pseudorandom functions (TPP-PRFs), including $\mathsf{XopEM}$, $\mathsf{SoEM22}$, $\mathsf{SUMPIP}$, and $\mathsf{DS\text{-}SoEM}$, partially answering Shi et al.’s open question. On the other hand, for block ciphers based on TPP-PRFs, we break the obstacle that this attack rely on online query by constructing decoupled XOR-type function, then propose an offline quantum attack on them that retains the tunable truncation parameter, $t$, a positive integer. Compared to previous results, our offline attack exhibits significantly reduced query complexity. Specifically, we reduce the number of queries to the encryption oracle from $\tilde{O}(2^{(n+t)/2})$ to $O(2^{t})$ with the same time complexity in the quantum query model, and enable its implementation in the classical query model, optimizing both the classical query complexity and time complexity from $\tilde{O}(2^{2n/3})$ to $\tilde{O}(2^{(2n-t)/3})$.
[…]
1 Introduction
Quantum computing, rooted in the principles of quantum mechanics, has potential speed-up advantages for solving specific classes of problems, such as combinatorial optimization [33, 35], image classification [32, 22], quantum architecture search  [28, 10], and so on [27, 19]. The rapid progress in quantum computing poses unprecedented threats to cryptographic security. Notably, Shor’s algorithm [25] and Grover’s algorithm [11] fundamentally threaten public-key and symmetric cryptography, respectively.
Simon’s algorithm [26] provides a crucial tool for quantum cryptanalysis by achieving exponential speedup over classical methods in finding the period of two-to-one periodic functions. This capability directly threatens the security of block ciphers (e.g., $3$-round Feistel [16], Even-Mansour cipher [17, 24, 34], etc.). In 2017, Grover-meets-Simon algorithm, which is presented by Leander and May [18], demonstrated that key whitening does not enhance security in the quantum-chosen-plaintext attack (q-CPA) setting on the FX construction. Subsequently, Grover-meets-Simon algorithm has played a pivotal role in researching the quantum security of symmetric cryptography, including $\mathsf{Feistel}$ [9, 14], $\mathsf{Lai\text{-}Massey}$ [20], $\mathsf{MISTY}$ structure [6], $\mathsf{SUM\text{-}ECBC}$ [13], and several Beyond-Birthday-Bound MACs [29].
The above attack relies on superposition queries to the encryption oracle (the Q2 model), which present significant implementation challenges. This has spurred growing interest in the attack that utilizes classical query and offline quantum computation (the Q1 model) in recent years. In 2019, Bonnetain et al. [4] proposed the offline Simon’s algorithm, dividing it into two distinct phases: the online phase for accessing the encryption oracle to prepare the superposition state and the offline computation phase for running the Grover iteration on this state to verify whether the function exhibits periodicity. They achieved key-recovery attacks on the FX and Even-Mansour constructions in the Q1 model, significantly optimizing attack complexities. In 2021, Bonnetain [3] explored the quantum implementation of the offline Simon’s algorithm. Further, Li et al. [19] proposed a quantum key recovery attack on $\mathsf{SoEM21}$ and $\mathsf{SoEMs1}$ in the Q1 model. Then, Sun et al. [31] systematically studied the quantum security of permutation-based pseudorandom functions, where the adversaries are restricted to classical queries and offline quantum computations.
As the National Institute of Standards and Technology points out, due to the limitations of quantum computer implementation technology, reducing the circuit depth of quantum attacks is essential for their practical implementation. Shi et al. [23] proposed a dedicated quantum attack on XOR-type function, which is written in the following form: $f(\alpha_{k},x)=g_{1}(x)\oplus g_{2}(\alpha_{k}\oplus x)$. In contrast to the parallelization of Grover-meets-Simon algorithm, their attack saved on circuit costs (e.g., the depth, width, and gate count), while requiring a quantum query complexity of $2^{(\kappa+t)/2}$ to the encryption oracle. Moreover, they analyzed that the attack has enabled improved cryptanalysis of several Beyond-Birthday-Bound (BBB) MACs, such as $\mathsf{SUM\text{-}ECBC}$, its variants $2\mathsf{K}$-$\mathsf{SUM\text{-}ECBC}$, $2\mathsf{K}$-$\mathsf{ECBC\_Plus}$, and $\mathsf{GCM\text{-}SIV2}$.
Motivations. Our work is motivated by two critical open problems arising from the preceding discussion.
First, the dedicated quantum attack proposed by Shi et al. [23] has only been applied to the BBB MACs mentioned above. A open question posed by the authors is to identify other cryptographic constructions that are vulnerable to this attack.
Second, this attack on these BBB MACs relies on superposition queries to the encryption oracle. This naturally leads to a more challenging question. Can the quantum attack on XOR-type function be performed in the Q1 model and other instances of block ciphers?
Our Contributions. Addressing these challenges of applicability and practicality, our work achieves the following breakthroughs:
  1. 1.
    we construct novel XOR-type function based on PolyMAC and two parallel permutation-based pseudorandom functions (TPP-PRFs). These constructions successfully provide practical instances of XOR-type function, thereby answering the open question posed by Shi et al. [23].
  2. 2.
    we propose the offline dedicated quantum attack on block ciphers based on TPP-PRFs by designing decoupled XOR-type function, e.g., $\mathsf{XopEM}$, $\mathsf{SoEM22}$, $\mathsf{SUMPIP}$, and $\mathsf{DS\text{-}SoEM}$ (See Table 1 for a comparison with previous and new quantum attacks on $\mathsf{SoEM22}$). Compared to the previous attack, this offline attack is reduced the query complexity from $\tilde{O}(2^{(n+t)/2})$ to $O(2^{t})$ in the Q2 model. And in the Q1 model, by retaining a tunable truncation parameter $t$, our work significantly optimizes the quantum-time/classical-query tradeoff from $\tilde{O}(2^{2n/3})$, as given by Sun et al. [31], to $\tilde{O}(2^{(2n-t)/3})$, where $t$ is a positive integer.
Organization. The rest of this paper is organized as follows: Section 2 provides the preliminaries, Section 3 discusses the offline dedicated quantum attacks on decoupled XOR-type function, and Section 4 provides a concise conclusion and discussion.
Table 1: Previous and new quantum attacks on $\mathsf{SoEM22}$
Algorithm Model Iteration Oracle query Time Query: c´ Reference
Grover-meets-Simon
algorithm [18] Q2 $2^{n/2}$ $O(2^{n/2}\cdot n)$ $O(2^{n/2}\cdot(n^{3}+T_{f}))$ $2n+\tau+1$ [24]
Dedicated attack
on XOR-type function [23] Q2 $2^{(n-t)/2}$ $O(2^{(n+t)/2}\cdot(n-t))$ $O(2^{(n-t)/2}\cdot(n^{3}+T_{f}))$ $2n-2t+\tau+1$ Ours
Offline dedicated attack
on $p$-XOR-type function Q2 $2^{(n-t)/2}$ $O(2^{t})$ $O(2^{(n-t)/2}\cdot(n^{3}+T_{p}))$ $2n-2t+\tau+1$ Ours
Offline Simon’s algorithm [4] Q1 $2^{2n/3}$ $O(2^{2n/3})$ $O(2^{2n/3}\cdot(n^{3}+T_{f}))$ $2n+\tau+1$ [31]
Offline dedicated attack
on $p$-XOR-type function Q1 $2^{(2n-t)/3}$ $O(2^{(2n-t)/3})$ $O(2^{(2n-t)/3}\cdot(n^{3}+T_{p}))$ $2n-2t+\tau+1$ Ours
  • 1
    The fourth column “Oracle query” indicates classical queries to the encryption oracle in the Q1 model and superposition queries in the Q2 model.
  • 2
    The fifth column “Time” represents the offline quantum computing time complexity in the Q1 model.
  • 3
    $T_{f}$ ($T_{p}$) is the time required to evaluate the function $f$ (the public function $p$) once.
2 Preliminaries
In this section, we review foundational algorithms, including Grover’s algorithm [11], Simon’s algorithm [26], the offline Simon’s algorithm [4], and the dedicated quantum attack on XOR-type function [23].
2.1 Grover’s algorithm
Grover’s algorithm [11] addresses the unstructured search problem, which involves finding a specific element that satisfies a given condition from a database of size $N=2^{n}$. Quantum Amplitude Amplification, as introduced in [5], represents a generalization of Grover’s algorithm.
Problem 1 (Grover’s problem).
Given a function $f:\{0,1\}^{n}\to\{0,1\}$ that can find the good element $x$ such that $f(x)=1$.
[…]
2.2 Simon’s algorithm
Simon’s problem involves finding the hidden period $s$ of a two-to-one periodic function. Simon’s algorithm [26] is designed to solve this problem and has an exponential speed advantage over classical algorithms.
Problem 2 (Simon’s Problem).
Given oracle access to $f:\{0,1\}^{n}\to\{0,1\}^{m}$, we promise that there exists a non-trivial $s\neq 0$ such that $f(x)=f(x\oplus s)$ for any $x$, find $s$.
Simon’s algorithm [26] works as follows: First, prepare a uniform superposition $|\phi_{1}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}|x\rangle|0\rangle$. Then, query the quantum oracle to obtain $f(x)$ in superposition $|\phi_{2}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}|x\rangle|f(x)\rangle$. And measure the second register. Let $f(z)$ be the measurement result. By omitting the second register, the resulting state is given by $|\phi_{3}\rangle=\frac{1}{\sqrt{2}}(|z\rangle\oplus|z\oplus s\rangle)$. Subsequently, apply Hadamard transformation to $|\phi_{3}\rangle$: $\frac{1}{\sqrt{2^{n+1}}}\sum_{y\in\{0,1\}^{n}}(-1)^{y\cdot z}(1+(-1)^{y\cdot s})|y\rangle$. Finally, measure the first register to obtain a random vector that is orthogonal to $s$. Repeating this process yields a system of linear equations, solved classically to extract $s$. Its key advantage is an exponential quantum speedup with only $O(n)$ queries, making it powerful for cryptanalysis of periodic functions in symmetric cryptography.
2.3 Grover-meets-Simon Algorithm
Grover-meets-Simon algorithm [18] represents a significant breakthrough in quantum cryptanalysis. It ingeniously combines Grover’s algorithm with Simon’s algorithm to solve key-recovery problems in symmetric ciphers.
Problem 3 (Grover-meets-Simon Problem).
Given oracle access to $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$, there exists a unique $k_{0}$ such that $f(k_{0},\cdot)$ hides a non-trivial period $s$, the goal is to find $k_{0}$ and the corresponding period $s$.
The algorithm’s framework is based on Grover’s search. It locates the correct key $k_{0}$ by performing a quantum search over the key space $\{0,1\}^{\kappa}$. In each Grover iteration, the algorithm utilizes Simon’s algorithm as a testing function to check whether $f(k,\cdot)$ is periodic or not. Given the output vectors $v_{1},\ldots,v_{c}$ from $c$ parallel Simon routines, it can be verified whether the rank of $\{v_{1},\ldots,v_{c}\}$ is at most n1n-1. Grover-meets-Simon algorithm, which requires $\mathcal{O}(n\cdot 2^{\kappa/2})$ quantum queries and $\mathcal{O}(n^{3}\cdot 2^{\kappa/2})$ time complexity, can be employed against various symmetric cryptographic schemes, such as the FX construction and some MACs.
2.4 The Offline Simon’s Algorithm
In 2019, Bonnetain et al. [4] presented the offline Simon’ algorithm for an asymmetric search of a period problem, with many cryptographic applications.
Problem 4 (Asymmetric Search of a Period Problem).
Let $F:\{0,1\}^{m}\times\{0,1\}^{n}\rightarrow\{0,1\}^{\ell}$ and $g:\{0,1\}^{n}\rightarrow\{0,1\}^{\ell}$ be two functions. We consider $F$ as a family of functions indexed by $\{0,1\}^{m}$ and write $F(i,\cdot)=f_{i}(\cdot)$. Assume that we are given quantum oracle access to $F$, and classical or quantum oracle access to $g$.
Assume that there exists exactly one $i_{0}\in\{0,1\}^{m}$ such that $f_{i}\oplus g$ has a hidden period, i.e., $\forall x\in\{0,1\}^{n}$, $f_{i_{0}}(x)\oplus g(x)=f_{i_{0}}(x\oplus s)\oplus g(x\oplus s)$ for some $s$. The goal is to find both $i_{0}$ and $s$.
The idea of Algorithm 1 is to decompose the process into two stages: the online phase prepares the superposition state $|\psi_{g}\rangle$ by querying the encryption oracle, and the offline quantum computation phase utilizes $|\psi_{g}\rangle$ to find the target solution via Grover iteration. Algorithm 2 illustrates how to implement the offline procedure, which checks whether the conditional period function $f\oplus g$ has a period without any query to $g$. The online phase is distinguished by the use of superposition queries to the encryption oracle in the Q2 model, whereas only classical queries are available in the Q1 model. The offline quantum computation phase runs in the same way.
Algorithm 1 The Offline Simon’s Algorithm  [4]
1:
[…]
3:
[…]
5:Start with $\left|{\psi_{g}}\right\rangle|0\rangle$.
6:Apply Hadamard transform to obtain
[…]
8:Apply $O(2^{m/2})$ Grover iterations to get
[…]
10:Note: The test oracle is a unitary operator that takes $|\psi_{g}\rangle|i\rangle$
11:as input and tests whether it $f_{i}\oplus g$ has a hidden period in superposition (see Algorithm 2 for details).
12:Measure the index $i\rightarrow i_{0}$
Algorithm 2 The procedure test that checks whether $f\oplus g$ has a period, without making any new queries to $g$ [4]
1:Start with $|\psi_{g}\rangle|b\rangle=\otimes^{cn}\left(\sum_{x\in\{0,1\}^{n}}|x\rangle|g(x)\rangle\right)|b\rangle$
2:Apply cncn $U_{f}$ to obtain
[…]
4:Apply $(H^{n}\otimes I_{m})^{cn}\otimes I_{1}$ to get
[…]
7:Compute $d:=\dim(\text{Span}(u_{1},\cdots,u_{cn}))$ Compute $d:=\dim(\mathrm{Span}(u_{1},\ldots,u_{cn}))$, set $r:=0$ if $d=n$ and $r:=1$ if $d<n$, and add $r$ to $b$. Then uncompute $d$ and $r$, and get
8: $\sum\limits_{\begin{subarray}{c}u_{1},\ldots,u_{cn}\\ x_{1},\ldots,x_{cn}\end{subarray}}(-1)^{u_{1}\cdot x_{1}}|{u_{1}}\rangle|(f\oplus g)(x_{1})\rangle\otimes\cdots\otimes(-1)^{u_{cn}\cdot x_{cn}}|{u_{cn}}\rangle|(f\oplus g)(x_{cn})\rangle\otimes|{b\oplus r}\rangle$.
9:Uncompute step 3-2 to get $|\psi_{g}\rangle|b\oplus r\rangle$
The complexity of Algorithm 1 is analyzed below.
Theorem 1 (Adapted from [4]).
Suppose that $m$ is in $O(n)$, Let $c$ be a sufficiently large constant. Consider Problem 4, there exists the index $i_{0}$ such that $f_{i_{0}}\oplus g$ has a period. Algorithm 1 identifies $i_{0}$ with probability $\Theta(1)$. In the Q2 model, it requires only $O(n)$ quantum queries to $g$ and $O(n2^{m/2})$ quantum queries to $F$. In the Q1 model, it requires $O(2^{n})$ classical queries to $g$ and $O(n2^{m/2})$ quantum queries to $F$. In both cases, the offline computation of Algorithm 1 is done in time $O((n^{3}+nT_{F})2^{m/2})$, where $T_{F}$ is the time needed to evaluate $F$ once.
We now review the specific estimates proposed by Bonnetain et al. [3] for the offline Simon’s algorithm.
Theorem 2 (Adapted from [3]).
Let $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$ satisfying Problem 3, $\tau$ be a positive integer and assume that $m\geq\log(4e(n+\kappa+\tau+1))$ and $\kappa\geq 7$, there exists a Grover-meets-Simon attack that uses $\pi/4\text{arcsin}\sqrt{2^{-\kappa}}$ iterations with $c$ quantum queries per iteration. The probability of the success rate being less than
$1-2^{-\tau}-\left(2^{-\tau/2-1}+2^{-\tau}+2^{-\kappa/2+1}\right)^{2}$ (1)
is lower than $2^{n+\kappa-\frac{2^{n}}{4c}}$, where $c=n+\kappa+\tau+1$.
2.5 Dedicated Quantum Attack on XOR-Type Function
XOR-type function constitutes a class of hidden periodic functions with a specific algebraic structure. Specifically, there exists a unique key $k_{0}$ such that the function $f(k_{0},\cdot)$ possesses a hidden period $s$, satisfying: $f(k_{0},x)=f(k_{0},x\oplus s)$. Furthermore, the function can be decomposed into two components: one independent of the key $k$, denoted as $g_{1}(x)$, and the other dependent on both the input $x$ and the key $k$, denoted as $g_{2}(k,x)$. This leads to the following expression for XOR-type function:
$f(k,x)=g_{1}(x)\oplus g_{2}(x\oplus\alpha_{k}),$
where $\alpha_{k}$ is an $n$-bit secret state derived from the key $k$.
The original space $\{0,1\}^{n}$ is partitioned into a $t$-dimensional linear subspace $\mathcal{L}=\{u:u=0^{n-t}\|*\}$ to construct a new function:
$F^{\mathcal{L}}(k,x^{l})=\bigoplus_{u\in\mathcal{L}}f(k,x^{l}\|0^{t}\oplus u).$
Using truncated technique, the parameters are decomposed as: $\alpha_{k}=\alpha_{k}^{l}\|\alpha_{k}^{r}$, $x=x^{l}\|x^{r}$, $s=s^{l}\|s^{r}$, $k=k^{l}\|k^{r}$, where $|\alpha_{k}^{l}|=|x^{l}|=|s^{l}|=|k^{l}|=n-t$ and $|\alpha_{k}^{r}|=|x^{r}|=|s^{r}|=|k^{r}|=t$. The resulting function $F^{\mathcal{L}}$ exhibits a truncated ntn-t-bit period $s^{l}$ when $\alpha_{k}^{l}=\alpha_{k_{0}}^{l}$, as stated in Property 1 [23].
Property 1 (Adapted from [23]).
Let $\mathcal{L}=\{u:u=0^{n-t}\|*\}$ and $f(k,x)$ be an XOR-type function such that $f(k,\cdot)$ has a non-trivial period $s$ when $k=k_{0}$. Then $F^{\mathcal{L}}(k,x^{l})=\bigoplus_{u\in\mathcal{L}}f(k,x^{l}\|0^{t}\oplus u)$ has an (nt)(n-t)-bit period $s^{l}$ for $\alpha^{l}_{k}=\alpha^{l}_{k_{0}}$.
In particular, if $f(k,x)=g_{1}(x)\oplus g_{2}(k\oplus x)$, then $F^{\mathcal{L}}(k^{l},x^{l})=\bigoplus_{u\in\mathcal{L}}f((k^{l}\oplus x^{l})\|0^{t}\oplus u)$ has an (nt)(n-t)-bit period $s^{l}$ for $k^{l}=k_{0}^{l}$.
The dedicated quantum attack on XOR-type function was proposed by Shi et al. [23], as presented in Theorem 3 and Algorithm 3.
Algorithm 3 Dedicated Attack on XOR-Type Function Using Truncated Technique [23]
Require: $\kappa,n,m,c^{\prime},t,r$, $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$ of XOR-type, to which we have quantum oracle access;
Ensure: the high (nt)(n-t) bits of $k_{0}$ and $s$, namely $k_{0}^{l}$ and $s^{l}$;
Prepare $(\kappa-t)+(n-t)c^{\prime}+mc^{\prime}$ qubits registers $|0^{\kappa-t}\rangle|0^{(n-t)c^{\prime}}\rangle|0^{mc^{\prime}}\rangle$.
1:Apply $(H^{\otimes\kappa-t}\otimes H^{\otimes(n-t)c^{\prime}}\otimes I_{2^{mc^{\prime}}})$ to obtain
[…]
3:Apply $U_{F\mathcal{L}}$ to get
[…]
5:Apply $I_{2^{\kappa-t}}\otimes H^{\otimes(n-t)c^{\prime}}\otimes I_{2^{mc^{\prime}}}$ to compute
6:
$|\psi\rangle=\sum\limits_{\begin{subarray}{c}k^{l}\in\{0,1\}^{\kappa-t},\\ x_{1}^{l},...,x_{c^{\prime}}^{l}\in\{0,1\}^{n-t},\\ v_{1},...,v_{c^{\prime}}\in\{0,1\}^{n-t}\end{subarray}}|k^{l}\rangle$ $(-1)^{x_{1}^{l}\cdot v_{1}}|v_{1}\rangle\ldots(-1)^{x_{c^{\prime}}^{l}\cdot v_{c^{\prime}}}|v_{c^{\prime}}\rangle$
$|F^{\mathcal{L}}(k^{l},x_{1}^{l})\rangle,\ldots,|F^{\mathcal{L}}(k^{l},x_{c^{\prime}}^{l})\rangle.$
7:Repeat Grover iteration $D_{|\psi\rangle}O_{test}$ for $r$ times to get $|\phi\rangle$:
[…]
9:Measure the first $((\kappa-t)+(n-t)c^{\prime})$-bit of $|\phi\rangle$ to get $k_{0}^{l}$ and $v_{1},\cdots,v_{c^{\prime}}$;
10:Compute the period $s^{l}$ by solving the equation system $s^{l}\cdot v_{i}=1$, $i=1,2,\ldots,c^{\prime}$;
Theorem 3.
(Adapted from [23]) Let $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$ be a function of XOR-type, then there exists a quantum attack that uses $2^{(\kappa-t)/2}$ iterations with $c^{\prime}$ quantum queries per iteration. The probability of falling below a success rate of
$1-2^{-\tau}-\left(2^{-\tau/2-1}+2^{-\tau}+2^{-(\kappa-t)/2+1}\right)^{2}$ (2)
is lower than $2^{n+\kappa-2t-\frac{2^{n-t}}{4c^{\prime}}}$, where $c^{\prime}=n+\kappa-2t+\tau+1$.
Through the internal parallelization strategy, the attack needs $2^{(\kappa-t)/2}$ Grover iterations, each requiring $c^{\prime}=2n-2t+\tau+1$ Simon subroutine queries and $2^{t}$ quantum queries to $f$, thus this attack requires $2^{(\kappa+t)/2}$ quantum queries to the encryption oracle. The high (nt)(n-t)-bit values $\alpha_{k_{0}}^{l}$ (or $k_{0}^{l}$) and the truncated period $s^{l}$ are recovered first. The remaining $t$-bit $k^{r}$ is then recovered using Grover-meets-Simon algorithm applied to the function $f^{\prime}(k^{r},x)=f^{\prime}(k^{r},x\oplus s)$, where $f^{\prime}(k^{r},x)=f(k^{l}_{0}\|k^{r},x)$. Note that this part requires only $2^{t/2}$ iterations. This approach demonstrates significant improvements over general parallelization in terms of circuit depth, width, and gate count.
3 Offline Dedicated Quantum Attack on $p$-XOR-Type Function
In this section, we systematically explore offline dedicated quantum attack on block ciphers based on $p$-XOR-type function. Beginning with a new instance based on XOR-type function for online attack, we illustrate the obstacle faced when converting online attacks into offline ones. Then, we present other instantiations of TPP-PRFs and provide an offline attack for them, as these instances can be constructed as decoupled XOR-type function, i.e., $p$-XOR-type function.
3.1 A New Instance for Online Attack and Obstacle to Converting to Offline Attack
In this subsection, we find a BBB MAC, called PolyMAC, which can be attacked using quantum queries in Ref. [23]. Subsequently, we analyze the obstacle that prevents this attack to offline setting.
3.1.1 Quantum attack on PolyMAC for online query
PolyMAC scheme [15] is a Double-block Hash-then-Sum construction based on polynomial evaluation in 2020. It uses two hashing keys $k_{1},k_{3}\in\{0,1\}^{n}$ and two encryption keys $k_{2},k_{4}\in\{0,1\}^{m}$. We consider the case with two-block messages as follows:
$\text{PolyMAC}(m_{1},m_{2})=E_{k_{2}}(k_{1}^{2}m_{1}\oplus k_{1}m_{2})\oplus E_{k_{4}}(k_{3}^{2}m_{1}\oplus k_{3}m_{2})$ (3)
where $M=m_{1}\|m_{2}$, and $|m_{1}|=|m_{2}|=n$.
Next, we briefly describe the construction of an XOR-type function based on PolyMAC. Let $\beta_{0}$, $\beta_{1}$ be two fixed strings in $\{0,1\}^{n}$, $\beta_{0}\neq\beta_{1}$, and $\alpha_{k}\in\{0,1\}^{n}$, we define the function as
$f:\{0,1\}^{n}\times\{0,1\}^{n}$ $\rightarrow\{0,1\}^{n}$ (4)
$(\alpha_{i},x)$ $\mapsto\text{PolyMAC}(\beta_{0},x)\oplus\text{PolyMAC}\left(\beta_{1},x\oplus\alpha_{i}\right).$
It is easy to verify that the period $s$ is $(k_{1}\oplus k_{3})(\beta_{0}\oplus\beta_{1})$ when $\alpha_{i}=k_{1}(\beta_{0}\oplus\beta_{1})$ or $\alpha_{i}=k_{3}(\beta_{0}\oplus\beta_{1})$. From Eq. (4), we observe that the secret state $\alpha_{i}$ is XORed with $x$ in the second term $\text{PolyMAC}(\beta_{1},x\oplus\alpha_{i})$ of $f(\alpha_{i},x)$, which exhibits the XOR property.
Let $\mathcal{L}=\{u:u=0^{n-t}\|*\}$, $\alpha_{i}=\alpha_{i}^{l}\|\alpha_{i}^{r}$, $x=x^{l}\|x^{r}$, $|\alpha_{i}^{l}|=|x^{l}|=n-t$ and $|\alpha_{i}^{r}|=|x^{r}|=t$. Now, we propose a new function $F^{\mathcal{L}}$ with truncated input.
$F^{\mathcal{L}}:\{0,1\}^{n-t}\times\{0,1\}^{n-t}$ $\rightarrow\{0,1\}^{m}$ (5)
$(\alpha_{i}^{l},x^{l})$ $\mapsto\bigoplus_{u\in\mathcal{L}}f(\alpha_{i}^{l},x^{l}\|0^{t}\oplus u).$
According to Algorithm 3, we can recover the high (nt)(n-t) bits of the period $s^{l}=(k_{1}\oplus k_{3})^{l}(\beta_{0}\oplus\beta_{1})^{l}$ when $\alpha_{i}^{l}=k_{1}^{l}(\beta_{0}\oplus\beta_{1})^{l}$ or $k_{3}^{l}(\beta_{0}\oplus\beta_{1})^{l}$ with truncated technique. Consider $\alpha_{i}^{l}=k_{1}^{l}(\beta_{0}\oplus\beta_{1})^{l}$ as an example. After that, we can define $f^{\prime}(\alpha_{i}^{r},x)=f(k_{1}^{l}(\beta_{0}\oplus\beta_{1})^{l}\|\alpha_{i}^{r},x)$. Obviously, $f^{\prime}(\alpha_{i}^{r},x)=f^{\prime}(\alpha_{i}^{r},x\oplus s)$ when $\alpha_{i}^{r}=k_{1}^{r}(\beta_{0}\oplus\beta_{1})^{r}$, it requires $O(2^{t/2})$ quantum queries to $f$.
3.1.2 Obstacle to Converting to Offline Attack
Although Section 3.1.1 identified a PolyMAC that can be attacked using dedicated attack proposed in Ref. [23], the difficulty of this attack lies in the requirement for quantum queries to $f$.
The obstacle that prevents the conversion of a dedicated quantum attack from an online to an offline setting is the tightly coupled property of XOR-type function $f(i,x)=g_{1}(x)\oplus g_{2}(x\oplus\alpha_{i})$ based on block ciphers. The core of the issue lies in the functional coupling between the two components, $g_{1}$ and $g_{2}$, which requires queries to the encryption oracle. This construction of $f(i,x)$ binds the entire query process to interactive online queries, thereby completely precluding any possibility of offline computation.
Specifically, consider the construction of XOR-type function based on PolyMAC given by Eq. (4), the first term $g_{1}(x)=\text{PolyMAC}(\beta_{0},x)$ and the second term $g_{2}(x\oplus\alpha_{i})=\text{PolyMAC}(\beta_{1},x\oplus\alpha_{i})$ of $f(\alpha_{i},x)$ both require quantum queries to the encryption oracle. We refer to the above structure as coupled XOR-type function. Similarly, XOR-type function constructed by several BBB MACs, which is proposed by Shi et al. [23], are all coupled XOR-type function.
Therefore, we introduce the decoupled XOR-type function to provide the offline attack.
decoupled XOR-type function. Let $f(i,x)$ be an XOR-type function in the following form $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\rightarrow\{0,1\}^{m}$, that is, $f(i,x)=g_{1}(x)\oplus g_{2}(x\oplus\alpha_{i})$, where $\alpha_{i}$ is the $n$-bit secret state determined by $i$. If the second term $g_{2}(x\oplus\alpha_{i})$ of $f(i,x)$ is a public function, which enables offline access, we call $f(i,x)$ is decoupled XOR-type function. To highlight its decoupled property, we denote it as XOR-type function with a public function ($p$-XOR-type function) and change its expression: $f(i,x)=g_{1}(x)\oplus p(x\oplus\alpha_{i})$, where $p(x\oplus\alpha_{i})$ represents a public function.
To facilitate discussion, we consider the case where $i$ is XORed on $x$, we introduce the following notation. Let $\mathcal{L}=\{u:u=0^{n-t}||*\}$, the new periodic function $F^{\mathcal{L}}(i^{l},x^{l})=\bigoplus\limits_{u\in\mathcal{L}}f(i^{l},x^{l}||0^{t}\oplus u)=G_{1}^{\mathcal{L}}(x^{l})\oplus P^{\mathcal{L}}(i^{l},x^{l})$, where $G_{1}^{\mathcal{L}}(x^{l})=\bigoplus\limits_{u\in\mathcal{L}}g_{1}(x^{l}\|0^{t}\oplus u)$, $P^{\mathcal{L}}$ consists of the public function $p(i,x)$, i.e., $P^{\mathcal{L}}(i^{l},x^{l})=\bigoplus\limits_{u\in\mathcal{L}}p((x^{l}\oplus i^{l})\|0^{t}\oplus u)$. It is obvious that $p$-XOR-type function is a special class of XOR-type function. Thus, Property 1 and Algorithm 3 proposed by Shi et al. [23] holds for $p$-XOR-type function.
3.2 Several Other Instances for Offline Attack
Given that the decoupling property of XOR-type function is crucial for enabling offline attack, we next explore a class of structure, TPP-PRFs, that inherently possess this characteristic, and present $p$-XOR-type function constructed from them.
3.2.1 Instantiation of TPP-PRFs
Let us show that $p$-XOR-type function constructed by instantiations of TPP-PRFs, and construct new periodic functions.
TPP-PRFs [7] TPP-PRFs with $n$-bit input $x$, which are constructed with two parallel permutations $P_{1}$ and $P_{2}$, present $2n/3$ bits of security in the classical setting, defined as follows.
$g:\{0,1\}^{n}$ $\rightarrow\{0,1\}^{n}$ (6)
$x$ $\mapsto l_{33}P_{1}(l_{13}(x)\oplus l_{14}(k_{1}))\oplus l_{34}P_{2}(l_{23}(x)\oplus l_{24}(k_{2}))\oplus e(x)\oplus C$
where $e(x)=l_{31}l_{11}(x)\oplus l_{32}l_{21}(x)$, the constant term $C=l_{31}l_{12}(k_{1})\oplus l_{32}l_{22}(k_{2})\oplus l_{35}(k_{3})\oplus l_{36}(k_{4})$, and $l_{ij}\neq 0$ for $i=1,2,3$, $j=3,4$.
Based on TPP-PRFs, we can construct a hidden periodic function that is a $p$-XOR-type function.
$f:\{0,1\}^{n}\times\{0,1\}^{n}$ $\rightarrow\{0,1\}^{n}$ (7)
$(i,x)$ $\mapsto g_{1}(x)\oplus p(i,x)$
$=l_{33}P_{1}(l_{13}(x\oplus l_{13}^{-1}l_{14}(k_{1}))\oplus l_{34}P_{2}(l_{23}(x\oplus l_{23}^{-1}l_{24}(k_{2}))$
$\quad\oplus l_{33}P_{1}\left(l_{13}(x)\right)\oplus l_{34}P_{2}\left(l_{23}(x\oplus i)\right)\oplus C,$
where $g_{1}(x)=g(x)\oplus l_{33}P_{1}\left(l_{13}(x)\right)\oplus e(x)$, and the public function $p:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\{0,1\}^{n}$, that is, $p(i,x)=l_{34}P_{2}\left(l_{23}(x\oplus i)\right)$. It can be verified that $f(i,x)=f(i,x\oplus l_{13}^{-1}l_{14}(k_{1}))$ when $i=l_{23}^{-1}l_{24}(k_{2})$ or $i=l_{23}^{-1}l_{24}(k_{2})\oplus l_{23}l_{13}^{-1}l_{14}(k_{1})$ in Ref. [31]. Since $i$ is XORed with $x$ in the public function $p(i,x)=l_{34}P_{2}(l_{23}(x\oplus i)$ of $f(i,x)$, the hidden periodic function $f(i,x)$ is a $p$-XOR-type function.
For $p$-XOR-type function constructed by TPP-PRFs described above, we can apply truncation techniques to obtain a new periodic function denoted as $F^{\mathcal{L}}$.
Let $\mathcal{L}=\{u:u=0^{n-t}\|*\}$, the (nt)(n-t)-bit key $k_{1}^{l}$ is decomposed into two parts: $k_{1}^{l_{1}}$ with $p$ bits and $k_{1}^{l_{2}}$ with (ntp)(n-t-p)-bits for $0\leq p\leq n-t$, where $t$ is a tunable positive integer truncation parameter. We define $G_{1}^{\mathcal{L}}:\{0,1\}^{p}\rightarrow\{0,1\}^{n}$ and the public function $P^{\mathcal{L}}:\{0,1\}^{2n-2t-p}\times\{0,1\}^{p}\rightarrow\{0,1\}^{n}$ as
$G_{1}^{\mathcal{L}}(x^{l_{1}})$ $=\bigoplus\limits_{u\in\mathcal{L}}g(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)\oplus e(x^{l_{1}}\|0^{n-t-p}),$
$P^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})$ $=\bigoplus\limits_{u\in\mathcal{L}}l_{33}P_{1}\left(l_{13}(x^{l_{1}}\|j^{l_{2}})\right)\oplus l_{34}P_{2}\left(l_{23}\left((x^{l_{1}}\|0^{n-t-p})\oplus i\right)\right).$
Now, we construct the new periodic function $F^{\mathcal{L}}:\{0,1\}^{2n-2t-p}\times\{0,1\}^{p}\rightarrow\{0,1\}^{n}$ as follows,
$F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=$ $G_{1}^{\mathcal{L}}(x^{l_{1}})\oplus P^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})$ (8)
$=$ $\bigoplus\limits_{u\in\mathcal{L}}g(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)\oplus l_{33}P_{1}\left(l_{13}(x^{l_{1}}\|j^{l_{2}})\right)$
$\oplus l_{34}P_{2}\left(l_{23}\left((x^{l_{1}}\|0^{n-t-p})\oplus i\right)\right)\oplus e(x^{l_{1}}\|0^{n-t-p}),$
where $l_{1}\in\{0,1\}^{p}$, $l_{2}\in\{0,1\}^{n-t-p}$, and $l=l_{1}\|l_{2}$.
In particular, this function has the period $s^{l_{1}}=l_{13}^{-1}l_{14}k_{1}^{l_{1}}$ if and only if $i^{l}\|j^{l_{2}}=l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$ (or $i^{l}\|j^{l_{2}}=l_{23}^{-1}l_{24}k_{2}^{l}\oplus l_{23}l_{13}^{-1}l_{14}k_{1}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$). As an example, we take $i^{l}\|j^{l_{2}}=l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$ to show that
$F^{\mathcal{L}}(l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}},x^{l_{1}})$
$\qquad=$ $\bigoplus\limits_{u\in\mathcal{L}}l_{33}P_{1}(l_{13}x^{l_{1}}\oplus l_{14}k_{1}^{l_{1}}\|l_{14}k_{1}^{l_{2}}\|0^{t}\oplus u)\oplus l_{34}P_{2}(l_{23}(x^{l_{1}}\|0^{n-t-p}\oplus l_{24}k_{2}^{l}\|0^{t}\oplus u)$
$\oplus l_{33}P_{1}(l_{13}x^{l_{1}}\|l_{14}k_{1}^{l_{2}}\|0^{t}\oplus u)\oplus l_{34}P_{2}(l_{23}(x^{l_{1}}\|0^{n-t-p}\oplus l_{23}^{-1}l_{24}k_{2}^{l})\|0^{t}\oplus u)\oplus C$
$\qquad=$ $\bigoplus\limits_{u\in\mathcal{L}}l_{33}P_{1}(l_{13}(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)\oplus l_{14}k_{1})\oplus l_{34}P_{2}(l_{23}(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)\oplus l_{24}k_{2})$
$\oplus l_{33}P_{1}\left(l_{13}(x^{l_{1}}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}\|0^{t}\oplus u)\right)\oplus l_{34}P_{2}\left(l_{23}(x^{l_{1}}\|0^{n-p-t})\oplus l_{24}k_{2}^{l}\|0^{t}\oplus u\right)\oplus C$
$\qquad=$ $F^{\mathcal{L}}(l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}},x^{l_{1}}\oplus l_{13}^{-1}l_{14}k_{1}^{l_{1}}).$
Next, we demonstrate that several instantiations of TPP-PRFs can construct $p$-XOR-type function, including $\mathsf{XopEM}$, $\mathsf{SoEM22}$, $\mathsf{SUMPIP}$, and $\mathsf{DS\text{-}SoEM}$.
The Xop construction [1] is defined as the bitwise XOR of the outputs from two distinct pseudorandom permutations (PRPs) applied to the same input $x$:
$\text{Xop}_{E_{1},E_{2}}(x)=E_{1}(x)\oplus E_{2}(x)$ (9)
where $E_{1}$ and $E_{2}$ represent the encryption algorithms of the respective PRPs.
The Xop construction is realized with two Even-Mansour ciphers, $EM_{1}(x)=P_{1}(x\oplus k_{1})\oplus k_{2}$ and $EM_{2}(x)=P_{2}(x\oplus k_{3})\oplus k_{4}$, resulting in the $\mathsf{XopEM}$ function.
$\mathsf{XopEM}(x)=P_{1}(x\oplus k_{1})\oplus P_{2}(x\oplus k_{3})\oplus k_{2}\oplus k_{4}$ (10)
We can obtain the $p$-XOR-type function $f(i,x)=P_{1}(x\oplus k_{1})\oplus P_{2}(x\oplus k_{3})\oplus P_{1}(x)\oplus P_{2}(x\oplus i)\oplus k_{2}\oplus k_{4}$, where $g_{1}(x)=\mathsf{XopEM}(x)$, a public function $p(i,x)=P_{1}(x)\oplus P_{2}(x\oplus i)$. This function has a period $s=k_{1}$ precisely when $i=k_{3}$.
Similarly, we define $G^{\mathcal{L}}:\{0,1\}^{p}\rightarrow\{0,1\}^{n}$ and $P^{\mathcal{L}}:\{0,1\}^{n-t}\times\{0,1\}^{n-t-p}\times\{0,1\}^{p}\rightarrow\{0,1\}^{n}$, where $l_{1}\in\{0,1\}^{p}$, $l_{2}\in\{0,1\}^{n-t-p}$, and $l=l_{1}\|l_{2}$:
$G^{\mathcal{L}}(x^{l_{1}})$ $=\bigoplus\limits_{u\in\mathcal{L}}\mathsf{XopEM}(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u),$
$P^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})$ $=\bigoplus\limits_{u\in\mathcal{L}}P_{1}(x^{l_{1}}\|j^{l_{2}}\|0^{t}\oplus u)\oplus P_{2}(((x^{l_{1}}\|0^{n-t-p})\oplus i^{l})\|0^{t}\oplus u).$
Consider a periodic function $F^{\mathcal{L}}:\{0,1\}^{n-t}\times\{0,1\}^{n-t-p}\times\{0,1\}^{p}\rightarrow\{0,1\}^{n}$,
$F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=G^{\mathcal{L}}(x^{l_{1}})\oplus P^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}}).$ (11)
It can be verified that $F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}}\oplus k_{1}^{l_{1}})$ when $i^{l}\|j^{l_{2}}=k_{3}^{l}\|k_{1}^{l_{2}}$ or $i^{l}\|j^{l_{2}}=k_{1}^{l}\oplus k_{3}^{l}\|k_{1}^{l_{2}}$.
SoEM22 [7]. The Sum of Even-Mansour (SoEM) construction employs two public permutations $P_{1}$, $P_{2}$, and two $n$-bit keys $k_{1}$, $k_{2}$. It represents a specific instance of the XOR construction where $k_{3}=k_{2}$ and $k_{4}=k_{1}$, defined as:
$\mathsf{SoEM22}(x)=P_{1}(x\oplus k_{1})\oplus P_{2}(x\oplus k_{2})\oplus k_{1}\oplus k_{2}.$ (12)
We can construct $p$-XOR-type function of the form $f(i,x)=\mathsf{SoEM22}(x)\oplus P_{1}(x)\oplus P_{2}(x\oplus i)$, where $P_{1}(x)\oplus P_{2}(x\oplus i)$ is a public function, which satisfies $f(i,x)=f(i,x\oplus k_{1})$ when $i=k_{2}$. The new periodic function is written as $F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=\bigoplus\limits_{u\in\mathcal{L}}\mathsf{SoEM22}(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)\oplus P_{1}(x^{l_{1}}\|j^{l_{2}}\|0^{t}\oplus u)\oplus P_{2}(((x^{l_{1}}\|0^{n-t-p})\oplus i^{l})\|0^{t}\oplus u)$, which has a period $k_{1}^{l_{1}}$ when $i^{l}\|j^{l_{2}}=k_{2}^{l}\|k_{1}^{l_{2}}$ or $i^{l}\|j^{l_{2}}=k_{1}^{l}\oplus k_{2}^{l}\|k_{1}^{l_{2}}$.
Consider $P_{1}=P$ and $P_{2}=P^{-1}$, the above $\mathsf{SoEM22}$ construction corresponds to $\mathsf{SUMPIP}$ [12]. In the same way, it can construct a $p$-XOR-type function $f(i,x)$ and generate a new periodic function $F^{\mathcal{L}}$ based on the truncation parameter.
DS-SoEM [2]. This construction is a domain-separated variant of the Xop construction, which is a sum of two Even-Mansour ciphers using a $d$-bit domain-separation constant. For an input $x\in\{0,1\}^{n-d}$, the encryption is defined as:
$\mathsf{DS\text{-}SoEM}(x)=$ $P\left(\left(x\oplus\mathrm{msb}_{n-d}(k_{1})\right)\|0^{d}\right)\oplus P\left(\left(x\oplus\mathrm{msb}_{n-d}(k_{2})\right)\|1^{d}\right)\oplus k_{1}\oplus k_{2},$ (13)
where $\mathrm{msb}_{n-d}$’ denotes truncation to the ndn-d most significant bits.
We define a $p$-XOR-type function $f(i,x)=\mathsf{DS\text{-}SoEM}(x)\oplus P(x\|0^{d})\oplus P\left((x\oplus i)\|1^{d}\right)$, which satisfies $f(i,x)=f\left(i,x\oplus\mathrm{msb}_{n-d}(k_{1})\right)$ when $i=\mathrm{msb}_{n-d}(k_{2})$, where $P(x\|0^{d})\oplus P\left((x\oplus i)\|1^{d}\right)$ is a public function. Then, we define the following function
$F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=$ $\bigoplus_{u\in\mathcal{L}}\mathsf{DS\text{-}SoEM}(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)\oplus P\left((x^{l_{1}}\|j^{l_{2}}\|0^{t}\oplus u)\|0^{d}\right)$ (14)
$\oplus P\left(\left((x^{l_{1}}\|0^{n-t-p}\oplus i^{l})\|0^{t}\oplus u\right)\|1^{d}\right).$
This function exhibits periodicity, that is, $F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}}\oplus[\mathrm{msb}_{n-d}(k_{1})]^{l_{1}})$ when $i^{l}\|j^{l_{2}}=[\mathrm{msb}_{n-d}(k_{2})]^{l}\|[\mathrm{msb}_{n-d}(k_{1})]^{l_{2}}$.
3.2.2 Offline Dedicated Quantum Attacks on Block Ciphers Based on TPP-PRFs
Based on the above $p$-XOR-type function constructed by TPP-PRFs, we now propose an offline dedicated quantum attack on them, utilizing the truncated technique in the Q1 (Q2) model, which reduces query complexity to the encryption oracle.
We present a new algorithm, termed the offline dedicated quantum attack on TPP-PRFs-based construction, which integrates the offline Simon’s algorithm with the dedicated quantum attack on $p$-XOR-type function. The main idea is to separate this attack into two phases: the online phase, which prepares the superposition state $|\psi_{G_{1}^{\mathcal{L}}}\rangle$ through classical or quantum queries to the encryption oracle $U_{g_{1}}$ (with classical queries requiring $2^{n}$ in the Q1 model by Algorithm 6, and quantum queries requiring parallel access to $2^{t}$ distinct $U_{g_{1}}$ in the Q2 model by Algorithm 7), and the offline computation phase, which performs the remaining quantum computations independently of the encryption oracle (with the test procedure checking whether the function $F^{\mathcal{L}}$ has a period without any new query to the encryption oracle in Algorithm 5). This separation significantly reduces the query complexity, as Grover iterations are executed offline after a limited number of online interactions.
Now, we introduce the offline dedicated quantum attack on block ciphers based on $p$-XOR-type function in the Q1 (Q2) model by Algorithm 4.
Algorithm 4 The Offline Dedicated Attack on block ciphers based on $p$-XOR-Type Function Using Truncated Technique
1:
2:Prepared $\left|\psi_{G_{1}^{\mathcal{L}}}\right\rangle=\otimes^{c^{\prime}}\left(\sum_{x^{l}\in\{0,1\}^{n-t}}|x^{l}\rangle|G_{1}^{\mathcal{L}}(x^{l})\rangle\right)$,
3:Let $\mathcal{L}=\{u:u=0^{n-t}||*\}$, $f(i,x)$ be a $p$-XOR-type function, and $f(i,\cdot)$ have a non-trivial period $s$ when $i=i_{0}$.
4:
5:The high (kt)(k-t)-bit of $i_{0}$ and (nt)(n-t)-bit of $s$, namely $i_{0}^{l}$ and $s^{l}$, s.t., $F^{\mathcal{L}}(i^{l},x^{l})=F^{\mathcal{L}}(i^{l},x^{l}\oplus s^{l})$ when $i^{l}=i^{l}_{0}$.
6:Start $\left|\psi_{G_{1}^{\mathcal{L}}}\right\rangle|0^{\kappa-t}\rangle$.
7:Apply $I_{c^{\prime}(n-t+m)}\otimes H^{\otimes\kappa-t}$ to obtain $|\Psi\rangle$:
[…]
9:Repeat $\mathcal{O}(2^{(\kappa-t)/2})$ Grover iterations:
[…]
11:Note that the test oracle is a unitary operator that takes $\left|\psi_{G_{1}^{\mathcal{L}}}\right\rangle|i^{l}\rangle$ as input, and tests whether $F^{\mathcal{L}}(i^{l},\cdot)$ has a hidden period (see Algorithm 5 for details).
12:Measure the index $i^{l}$ to obtain $i_{0}^{l}$.
Algorithm 5 The procedure checks whether a function $F^{\mathcal{L}}(i^{l},x^{l})=G_{1}^{\mathcal{L}}(x^{l})\oplus P^{\mathcal{L}}(i^{l},x^{l})$ has a period, without making any new queries to $g_{1}$.
1:
2:Prepared $|\psi_{G_{1}^{\mathcal{L}}}\rangle$.
3:Start $|\psi_{G_{1}^{\mathcal{L}}}\rangle|b\rangle$:
[…]
5:Apply $c^{\prime}$ $U_{P^{\mathcal{L}}}$ to obtain $|\psi_{F^{\mathcal{L}}}\rangle$:
6:
$|\psi_{F^{\mathcal{L}}}\rangle=$ $\otimes^{c^{\prime}}\left(\sum_{x^{l}\in\{0,1\}^{n-t}}|x^{l}\rangle|G_{1}^{\mathcal{L}}(x^{l})\oplus P^{\mathcal{L}}(x^{l})\rangle\right)|b\rangle$
$=$ $\otimes^{c^{\prime}}\left(\sum_{x^{l}\in\{0,1\}^{n-t}}|x^{l}\rangle|F^{\mathcal{L}}(x^{l})\rangle\right)|b\rangle.$
7:Apply $(H^{\otimes(n-t)}\otimes I_{m})^{\otimes c^{\prime}}\otimes I_{1}$ to $|\psi_{F^{\mathcal{L}}}\rangle$, to get:
8:
$\left(\sum\limits_{v_{1},x_{1}^{l}\in\{0,1\}^{n-t}}(-1)^{v_{1}\cdot x_{1}^{l}}|v_{1}\rangle|F^{\mathcal{L}}(x_{1}^{l})\rangle\right)\otimes\cdots$
$\otimes\left(\sum\limits_{v_{c^{\prime}},x_{c^{\prime}}^{l}\in\{0,1\}^{n-t}}(-1)^{v_{c^{\prime}}\cdot x_{c^{\prime}}^{l}}|v_{c^{\prime}}\rangle|F^{\mathcal{L}}(x_{c^{\prime}}^{l})\rangle\right)|b\rangle.$
9:Compute $d:=\dim(\text{Span}(v_{1},\dots,v_{c^{\prime}}))$.
10:if $d=n$ then
11:  set $r:=0$
12:else
13:  set $r:=1$, and add $r$ to $b$
14:end if
15:Uncompute steps 3-2 to obtain:
[…]
Algorithm 6 Prepare $|\psi_{G_{1}^{\mathcal{L}}}\rangle$ in the Q1 model
1:
2:Classical query to the encryption oracle $g_{1}$.
[…]
4:Apply $(H^{\otimes(n-t)}\otimes I_{m})^{\otimes c^{\prime}}$ to obtain:
[…]
6:For each $x^{l}\in\{0,1\}^{n-t}$, classical queries to $g_{1}$ to get $\bigoplus\limits_{u\in\mathcal{L}}g_{1}(x^{l}||0^{t}\oplus u)$.
7:Apply unitary operation to obtain:
8:
$|\psi_{G_{1}^{\mathcal{L}}}\rangle$ $=\otimes^{c^{\prime}}\left(\sum_{x^{l}\in\{0,1\}^{n-t}}|x^{l}\rangle|\bigoplus\limits_{u\in\mathcal{L}}g_{1}(x^{l}||0^{t}\oplus u)\rangle\right)$
$=\otimes^{c^{\prime}}\left(\sum_{x^{l}\in\{0,1\}^{n-t}}|x^{l}\rangle|G_{1}^{\mathcal{L}}(x^{l})\rangle\right).$
Algorithm 7 Prepare $|\psi_{G_{1}^{\mathcal{L}}}\rangle$ in the Q2 model
1:
2:quantum queries to the encryption oracle $g_{1}$.
[…]
4:Apply $(H^{\otimes(n-t)}\otimes I_{m})^{\otimes c^{\prime}}$ to obtain:
[…]
6:The XOR sum of $2^{t}$ distinct $U_{g_{1}}$, which can be implemented by using CNOTs, we can obtain $|\psi_{G_{1}^{\mathcal{L}}}\rangle$:
7:
$|\psi_{G_{1}^{\mathcal{L}}}\rangle$ $=\otimes^{c^{\prime}}\left(\sum_{x^{l}\in\{0,1\}^{n-t}}|x^{l}\rangle|G_{1}^{\mathcal{L}}(x^{l})\rangle\right).$
The overall quantum attack aims to find the high $(\kappa-t)$ bits $i_{0}^{l}$ such that it $F^{\mathcal{L}}(i_{0}^{l},\cdot)$ exhibits a non-trivial period $s^{l}$. Starting with the superposition state $\left|\psi_{G_{1}^{\mathcal{L}}}\right\rangle=\otimes^{c^{\prime}}\left(\sum_{x^{l}\in\{0,1\}^{n-t}}|x^{l}\rangle|G_{1}^{\mathcal{L}}(x^{l})\rangle\right)$, the input is $\left|\psi_{G_{1}^{\mathcal{L}}}\right\rangle|0^{\kappa-t}\rangle$ of $(\kappa-t)+c^{\prime}(n-t)+c^{\prime}m$ qubits. The initial state $|\Psi\rangle$ required for the Grover iteration is obtained by several $H$ transforms. The iterative function is $G=D_{|\Psi\rangle}O_{\text{test}}$ for $\mathcal{O}(2^{(\kappa-t)/2})$ iterations, where $D_{|\Psi\rangle}=(2|\Psi\rangle\langle\Psi|-I)$. The final step is measuring the index $i^{l}$ to obtain $i_{0}^{l}$. The implementation of $U_{P^{\mathcal{L}}}$ and $O_{\text{test}}$ in Algorithm 4 follows an approach similar to that used in Ref. [23]. We briefly illustrate the implementation of $U_{P^{\mathcal{L}}}$. The function
$P^{\mathcal{L}}(i^{l},x^{l})=\bigoplus_{u\in\mathcal{L}}p((x^{l}\oplus i^{l})\|0^{t}\oplus u)$
is the XOR sum of $2^{t}$ public functions $p$, which can be implemented in parallel without a sequential relationship, where $t$ is a tunable truncation parameter.
To get a usable success probability after $r$ iterations, we need $c^{\prime}=n+\kappa-2t+\pi+1$ queries per iteration. Next, we give the complexity and success probability of the offline dedicated quantum attack on block ciphers based on $p$-XOR-type function in the Q1 (Q2) model by Theorem 4.
Theorem 4.
Let $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$ be a $p$-XOR-type function, then there exists the offline dedicated quantum attack on $f$ by Algorithm 4. This attack can recover $i_{0}^{l}\in\{0,1\}^{\kappa-t}$ with a time complexity of $O(2^{(\kappa-t)/2}\cdot(n^{3}+T_{p}))$, requiring $O(2^{n})$ classical queries to the encryption oracle in the Q1 model, and $O(2^{t})$ quantum queries in the Q2 model. The probability of falling below a success rate of
$1-2^{-\tau}-\left(2^{-\tau/2-1}+2^{-\tau}+2^{-(\kappa-t)/2+1}\right)^{2}$ (15)
is lower than $2^{n+\kappa-2t-\frac{2^{n-t}}{4c^{\prime}}}$, where $c^{\prime}=n+\kappa-2t+\tau+1$ and $T_{p}$ is the time required to compute the public function $p$ once.
For $p$-XOR-type functions and new periodic functions $F^{\mathcal{L}}$ constructed from some instantiations of TPP-PRFs in Section 3.2.1, we can implement offline dedicated attacks on them to recover the period of $F^{\mathcal{L}}$ by Theorem 4.
In the Q1 model, this attack on block ciphers based on TPP-PRFs requires $O(2^{p+t})$ classical queries to the encryption oracle and $O(2^{(2n-2t-p)/2}\cdot n^{3})$ offline quantum computation time. Note that each iteration of $p$ can be completed in $O(1)$ time. Let $D$ and $T$ be the number of classical queries and the time of offline quantum computations. A classical attack exhibits a quantum-time/classical-query tradeoff of $T^{2}D=2^{2n-t}$, which is balanced at $T=D=2^{(2n-t)/3}$ for a tunable truncation parameter $t$.
The remaining keys can be recovered as follows. Taking TPP-PRFs as an example, after we obtain $l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$, Simon’s algorithm is used to recover $l_{13}^{-1}l_{14}k_{1}^{l_{1}}$ by requiring $O(2^{p+t})$ classical queries, and its offline computation requires $O(n^{3})$ time. Then, we can easily recover the remaining $t$-bit by applying the offline Simon’s algorithm on $f^{\prime}(k^{r},x)=f(k_{2}^{l}\|i^{r},x)$, which has the period $s$ when $i^{r}=k_{0}^{r}$.
In the Q2 model, consider the offline attack proposed in Theorem 4, the construction of $F^{\mathcal{L}}$ differs from that in Section 3.2.1 in that there is unnecessary to partition the (nt)(n-t)-bit subkey $k_{1}^{l}$ into $k_{1}^{l_{1}}$ with $p$-bit and $k_{1}^{l_{2}}$ with (ntp)(n-t-p)-bit.
Taking the $\mathsf{SoEM22}$ construction as an example, we can construct $p$-XOR-type function $f(i,x)=\mathsf{SoEM22}\oplus P_{1}(x)\oplus P_{2}(x\oplus i)$ with the public function $P_{2}(x\oplus i)$, then obtain a new periodic function $F^{\mathcal{L}}(i^{l},x^{l})=\bigoplus\limits_{u\in\mathcal{L}}\mathsf{SoEM22}(x^{l}\|0^{t}\oplus u)\oplus P_{1}(x^{l}\|0^{t}\oplus u)\oplus P_{2}\left((x^{l}\oplus i^{l})\|0^{t}\oplus u\right)$ that $F^{\mathcal{L}}(i^{l},x^{l})$ has an (nt)(n-t)-bit period $s^{l}=k^{l}_{1}$ when $i^{l}=k^{l}_{2}$. Based on this, we can prepare the state $|\psi_{G_{1}^{\mathcal{L}}}\rangle$ by requiring $2^{t}$ quantum queries to $\mathsf{SoEM22}$ by Algorithm 7. Then, we apply the offline attack by Algorithm 4, which successfully recovers the period $k^{l}_{1}$ with $O\left(2^{(n-t)/2}\cdot(n^{3}+T_{p})\right)$ offline quantum computation time. Compared to the attack by Shi et al. [23], our method takes the same time complexity but reduces the quantum query complexity from $O\left(2^{(n+t)/2}\cdot(n-t)\right)$ to $O(2^{t})$, demonstrating the advantage of our offline attack in the Q2 model.
4 Conclusion
This work successfully analyzed the quantum security of more instances of XOR-type function based on PolyMAC and TPP-PRFs, thereby resolving the open question raised by Shi et al. [23]. The dedicated quantum attack on BBB MACs required the adversary to perform superposition queries. We introduce a decoupled XOR-type function, termed $p$-XOR-type function, and propose an offline dedicated quantum attack on block ciphers based on TPP-PRFs. This attack yields significant efficiency improvements. In the Q2 model, the complexity of quantum query is reduced from $O(2^{(n+t)/2}\cdot(n-t))$ to $O(2^{t})$ with the same time complexity. In the Q1 model, the offline attack can recover the key of TPP-PRFs, i.e., $l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$, by making $O(2^{p+t})$ classical queries and performing $O(2^{(2n-2t-p)/2}\cdot n^{3})$ offline computation time. It features a quantum-time/classical-query tradeoff of $T^{2}D=2^{2n-t}$, requiring $T=D=O(2^{(2n-t)/3})$, where $t$ is a tunable truncation parameter. Finally, we provide new insights into the quantum security of cryptographic schemes, including TPP-PRFs and specific instantiations such as $\mathsf{XopEM}$, $\mathsf{SoEM22}$, $\mathsf{SUMPIP}$, and $\mathsf{DS\text{-}SoEM}$.
Our results open several promising avenues for future research: First, explore the quantum security of more cryptographic structures that can construct decoupled XOR-type function in the Q1 model. Second, develop more generic and effective offline dedicated quantum attacks by identifying algebraic properties that apply to assess the quantum security of block ciphers, including $\mathsf{Feistel}$ [9, 14, 30], $\mathsf{MISTY}$ [6], $\mathsf{Lai\text{-}Massey}$ structure [20], and other permutation-based pseudorandom functions (e.g., $\mathsf{EDM}$ [8] and $\mathsf{EDMD}$ structure [21]).
Acknowledgments
This work is supported by National Natural Science Foundation of China (Grant Nos. 62372048, 62272056, 62371069 )