### v3

# Offline Dedicated Quantum Attacks on Block Cipher Constructions Based on Two Parallel Permutation-Based Pseudorandom Functions

Abstract
Quantum cryptanalysis is essential for evaluating the security of cryptographic systems against the threat of quantum computing. Recently, Shi et al. introduced a dedicated quantum attack on block cipher constructions based on XOR-type functions, which greatly reduces the required resources (including circuit depth, width, and the number of gates) compared to the parallel Grover-meets-Simon algorithm. Here, our contribution is in two aspects. On the one hand, we discover new cryptographic structures amenable to this attack: PolyMAC and constructions based on two parallel permutation-based pseudorandom functions (TPP-PRFs), including $\mathsf{XopEM}$, $\mathsf{SoEM22}$, $\mathsf{SUMPIP}$, and $\mathsf{DS\text{-}SoEM}$, thereby answering Shi et al.’s open question. On the other hand, for constructions based on TPP-PRFs, we break the obstacle that this attack relies on online query by constructing decoupled XOR-type functions, then propose an offline quantum attack on them. Compared to previous results, our offline attack exhibits significantly reduced query complexity. Specifically, the number of queries to the encryption oracle is reduced from $O(2^{(n+t)/2}\cdot(n-t))$ to $O(2^{t}\cdot(n-t))$ in the quantum query model, where $0<t<n$, $t$ is a truncation parameter, and $n$ is the input length of constructions. Further, we enable its implementation in the classical query model, optimizing both the classical query complexity and time complexity from $\tilde{O}(2^{2n/3})$ to $\tilde{O}(2^{(2n-t)/3})$.
publicationid: pubid:
I Introduction
Quantum computing, rooted in the principles of quantum mechanics, exhibits speed-up advantages for specific computational problems, such as linear system solving [harrow2009quantum, wan2018asymptotic], quantum simulation [Li2025quantumsimulation], and machine learning [li2025efficient, huang2021power]. As a special kind of quantum algorithm for Noisy Intermediate Scale Quantum devices, the Variational Quantum Algorithm (VQA) [cerezo2021variational, wang2021noise, su2025topology] is currently being extensively studied, showing potential advantages in tasks such as combinatorial optimization [wu2025resource, wu2025], image classification [wu2024quantum, sein2025image], and so on [peruzzo2014variational, song2024quantum].
The development of quantum computing poses unprecedented threats to cryptographic security. Notably, Shor’s algorithm [shor1994algorithms] and Grover’s algorithm [grover1996fast] fundamentally threaten public-key and symmetric cryptography, respectively. Meanwhile, VQAs have also garnered attention in quantum cryptanalysis [wang2025reducing, wang2022variational], particularly for analyzing hard problems underlying post-quantum cryptography, such as Shortest Vector Problem [albrecht2023variational], Learning With Errors Problem [zheng2025quantum, zeng2025analysis].
Simon’s algorithm [simon1997power] provides a crucial tool for quantum cryptanalysis by achieving exponential speedup over classical methods in finding the period of two-to-one periodic functions. This capability directly threatens the security of block ciphers (e.g., $3$-round Feistel [kuwakado2010quantum], Even-Mansour cipher [kuwakado2012security, shinagawa2022quantum, zhang2023quantum], etc.). In 2017, Grover-meets-Simon algorithm, which is presented by Leander and May [leander2017grover], demonstrated that key whitening does not enhance security in the quantum-chosen-plaintext attack (q-CPA) setting on the FX construction. Subsequently, Grover-meets-Simon algorithm has played a pivotal role in researching the quantum security of symmetric cryptography, including $\mathsf{Feistel}$ [dong2018quantum, ito2019quantum], $\mathsf{Lai\text{-}Massey}$ [mao2022quantum], $\mathsf{MISTY}$ structure [canale2022simon], $\mathsf{SUM\text{-}ECBC}$ [guo2021attacks], and several Beyond-Birthday-Bound MACs [sun2023quantum].
The above attacks, based on Simon’s and Grover-meets-Simon algorithms, rely on superposition queries to the encryption oracle (the Q2 model), presenting significant implementation challenges. This has spurred growing interest in a quantum attack that utilizes classical queries and offline quantum computation (the Q1 model) in recent years. In 2019, Bonnetain et al. [bonnetain2019quantum] proposed the offline Simon’s algorithm and achieved key-recovery attacks on the FX and Even-Mansour constructions in the Q1 model. Following this, Li et al. [li2025quantum] proposed a quantum key recovery attack on $\mathsf{SoEM21}$ and $\mathsf{SoEMs1}$ in the Q1 model. Further, Sun et al. [sun2025quantum] systematically studied the quantum security of block cipher constructions based on permutation-based pseudorandom functions in the Q1 model.
Due to the limitations of quantum computer implementation technology, reducing the circuit depth of quantum attacks is essential for the practical implementation. Shi et al. [shi2024dedicated] proposed a dedicated quantum attack on constructions based on XOR-type functions of the form $f(\alpha_{k},x)=g_{1}(x)\oplus g_{2}(\alpha_{k}\oplus x)$. In contrast to previous attacks, their attack reduced circuit costs, including depth, width, and gate count. Moreover, they applied this attack to several Beyond-Birthday-Bound (BBB) MACs, such as $\mathsf{SUM\text{-}ECBC}$, its variants $2\mathsf{K}$-$\mathsf{SUM\text{-}ECBC}$, $2\mathsf{K}$-$\mathsf{ECBC\_Plus}$, and $\mathsf{GCM\text{-}SIV2}$. For $\mathsf{SUM\text{-}ECBC}$, it simultaneously achieved savings of at least $62.3\%$ in depth, $19.5\%$ in width, and $22.2\%$ in gate count.
Motivations. Our work is motivated by two critical questions arising from the above dedicated attack.
First, the dedicated quantum attack proposed by Shi et al. [shi2024dedicated] is currently limited to the BBB MACs mentioned above. They posed an open question that is to identify other cryptographic constructions that are vulnerable to this attack.
Second, this attack on these BBB MACs relies on quantum queries to the encryption oracle. This naturally leads to a more challenging question, that is, can this quantum attack be performed in the offline setting?
Our Contributions. Addressing these challenges of applicability and practicality, our work achieves the following results:
  1. 1.
    We construct XOR-type functions based on PolyMAC and two parallel permutation-based pseudorandom functions (TPP-PRFs), that is, we identify new instances that are vulnerable to the attack presented by Shi et al. [shi2024dedicated], thereby answering their open question.
  2. 2.
    We propose an OFFLINE dedicated quantum attack on constructions based on TPP-PRFs by designing decoupled XOR-type functions, successfully applying it to constructions like $\mathsf{XopEM}$, $\mathsf{SoEM22}$, $\mathsf{SUMPIP}$, and $\mathsf{DS\text{-}SoEM}$. Compared to the previous attack, the query complexity in our attack is reduced from $O(2^{(n+t)/2}\cdot(n-t))$ to $O(2^{t}\cdot(n-t))$ in the Q2 model, where $0<t<n$, $t$ is a truncation parameter, and $n$ is the input length of constructions. Furthermore, we significantly optimize the time complexity and classical query complexity from $\tilde{O}(2^{2n/3})$, as given by Sun et al. [sun2025quantum], to $\tilde{O}(2^{(2n-t)/3})$ in the Q1 model.
Organization. The rest of this paper is organized as follows: Section II provides the preliminaries, Section III discusses an offline dedicated quantum attack on block ciphers by constructing decoupled XOR-type functions, and Section IV provides a concise conclusion and discussion.
II Preliminaries
In this section, we review foundational algorithms, including Grover’s algorithm [grover1996fast], Simon’s algorithm [simon1997power], the offline Simon’s algorithm [bonnetain2019quantum], and the dedicated quantum attack on constructions based on XOR-type functions [shi2024dedicated].
II-A Grover’s algorithm
Grover’s algorithm [grover1996fast] addresses the unstructured search problem, which involves finding a specific element that satisfies a given condition from a database of size $N=2^{n}$. Quantum Amplitude Amplification, introduced in [brassard2000quantum], is a generalization of Grover’s algorithm.
Problem 1 (Grover’s problem)
Given a function $f:\{0,1\}^{n}\to\{0,1\}$ that can find the good element $x$ such that $f(x)=1$.
[…]
II-B Simon’s algorithm
Simon’s problem involves finding the unique hidden period $s$ of a two-to-one periodic function. Simon’s algorithm [simon1997power] solves this problem and provides an exponential speed advantage over classical algorithms.
Problem 2 (Simon’s Problem)
Given oracle access to a function $f:\{0,1\}^{n}\to\{0,1\}^{m}$, we are promised that there exists a non-trivial period $s\neq 0$ such that $f(x)=f(x\oplus s)$ for any $x\in\{0,1\}^{n}$. The goal is to find $s$.
Simon’s algorithm [simon1997power] works as follows: First, prepare a uniform superposition state $|\phi_{1}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}|x\rangle|0\rangle$. Then, query the quantum oracle to obtain $f(x)$ in superposition $|\phi_{2}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}|x\rangle|f(x)\rangle$. Then, measure the second register. Let the measurement outcome be $f(z)$. By tracing out the second register, the first register collapses to a superposition $|\phi_{3}\rangle=\frac{1}{\sqrt{2}}(|z\rangle+|z\oplus s\rangle)$. Subsequently, apply the Hadamard transformation to $|\phi_{3}\rangle$: $\frac{1}{\sqrt{2^{n+1}}}\sum_{y\in\{0,1\}^{n}}(-1)^{y\cdot z}(1+(-1)^{y\cdot s})|y\rangle$. Finally, measure the first register to obtain a random vector $y$ such that $y\cdot s=0\mod 2$. Repeating this process yields a system of linear equations, which is then solved classically to efficiently extract $s$. The key advantage of Simon’s algorithm is an exponential quantum speedup with only $O(n)$ queries, making it a powerful tool for the cryptanalysis of periodic functions in symmetric cryptography.
II-C The Offline Simon’s Algorithm
In 2019, Bonnetain et al. [bonnetain2019quantum] presented the offline Simon’s algorithm to solve the asymmetric search of a period problem, which has some cryptographic applications.
Problem 3 (Asymmetric Search of a Period Problem)
Let $F:\{0,1\}^{m}\times\{0,1\}^{n}\rightarrow\{0,1\}^{\ell}$ and $g:\{0,1\}^{n}\rightarrow\{0,1\}^{\ell}$ be two functions. We consider $F$ as a family of functions indexed by $\{0,1\}^{m}$ and write $F(i,\cdot)=f_{i}(\cdot)$. Assume that we are given quantum oracle access to $F$, and classical or quantum oracle access to $g$.
Assume that there exists exactly one $i_{0}\in\{0,1\}^{m}$ such that $f_{i}\oplus g$ has a hidden period, i.e., $\forall x\in\{0,1\}^{n}$, $f_{i_{0}}(x)\oplus g(x)=f_{i_{0}}(x\oplus s)\oplus g(x\oplus s)$ for some $s$. The goal is to find both $i_{0}$ and $s$.
Algorithm 1 The Offline Simon’s Algorithm  [bonnetain2019quantum]
1:
[…]
3:
[…]
5:Start with $\left|{\psi_{g}}\right\rangle|0\rangle$.
6:Apply Hadamard transform to obtain
[…]
8:Apply $O(2^{m/2})$ Grover iterations to get
[…]
10:Note: The test oracle is a unitary operator that takes $|\psi_{g}\rangle|i\rangle$
11:as input and tests whether it $f_{i}\oplus g$ has a hidden period in superposition (see Algorithm 2 for details).
12:Measure the index $i\rightarrow i_{0}$
Algorithm 2 The procedure test that checks whether $f\oplus g$ has a period, without making any new queries to $g$ [bonnetain2019quantum]
1:Start with $|\psi_{g}\rangle|b\rangle=\otimes^{cn}\left(\sum_{x\in\{0,1\}^{n}}|x\rangle|g(x)\rangle\right)|b\rangle$
2:Apply cncn $U_{f}$ to obtain
[…]
4:Apply $(H^{n}\otimes I_{m})^{cn}\otimes I_{1}$ to get
[…]
7:Compute $d:=\dim(\text{Span}(u_{1},\cdots,u_{cn}))$ Compute $d:=\dim(\mathrm{Span}(u_{1},\ldots,u_{cn}))$, set $r:=0$ if $d=n$ and $r:=1$ if $d<n$, and add $r$ to $b$. Then uncompute $d$ and $r$, and get
8: $\sum\limits_{\begin{subarray}{c}u_{1},\ldots,u_{cn}\\ x_{1},\ldots,x_{cn}\end{subarray}}(-1)^{u_{1}\cdot x_{1}}|{u_{1}}\rangle|(f\oplus g)(x_{1})\rangle\otimes\cdots\otimes(-1)^{u_{cn}\cdot x_{cn}}|{u_{cn}}\rangle|(f\oplus g)(x_{cn})\rangle\otimes|{b\oplus r}\rangle$.
9:Uncompute step 3-2 to get $|\psi_{g}\rangle|b\oplus r\rangle$
The idea of Algorithm 1 is to decompose the process into two stages: the online phase prepares the superposition state $|\psi_{g}\rangle$ by querying the encryption oracle, and the offline quantum computation phase utilizes $|\psi_{g}\rangle$ to find the target solution using the Simon procedure. Algorithm 2 illustrates how to implement the offline procedure, which checks whether the conditional period function $f\oplus g$ has a period without any query to $g$. The Q1 and Q2 models fundamentally differ only in the online phase. The superposition queries to the encryption oracle are permitted in the Q2 model, whereas the classical queries are restricted in the Q1 model. The subsequent offline quantum computation phase remains identical in both models.
The complexity of Algorithm 1 is analyzed below.
Theorem 1
(Adapted from [bonnetain2019quantum]) Assume that $m=O(n)$, and let $c$ be a sufficiently large constant. Consider Problem 3, there exists a unique index $i_{0}$ such that $f_{i_{0}}\oplus g$ has a period. Algorithm 1 identifies $i_{0}$ with probability $\Theta(1)$. In the Q2 model, it requires only $O(n)$ quantum queries to $g$ and $O(n2^{m/2})$ quantum queries to $F$. In the Q1 model, it requires $O(2^{n})$ classical queries to $g$ and $O(n2^{m/2})$ quantum queries to $F$. In both cases, the offline computation time of Algorithm 1 is $O((n^{3}+nT_{F})2^{m/2})$, where $T_{F}$ is the time needed to evaluate $F$ once.
We now review the specific estimates proposed by Bonnetain [bonnetain2021tight] for the offline Simon’s algorithm.
Theorem 2
(Adapted from [bonnetain2021tight]) Let $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$, and assume that there exists a unique $k_{0}$ such that $f(k_{0},x)=f(k_{0},x\oplus s)$, $\tau$ be a positive integer, $m\geq\log(4e(n+\kappa+\tau+1))$, and $\kappa\geq 7$. There exists the offline Simon attack that uses $\pi/4\text{arcsin}\sqrt{2^{-\kappa}}$ iterations with $c$ quantum queries per iteration. The probability of the success rate being less than
$1-2^{-\tau}-\left(2^{-\tau/2-1}+2^{-\tau}+2^{-\kappa/2+1}\right)^{2}$ (1)
is lower than $2^{n+\kappa-\frac{2^{n}}{4c}}$, where $c=n+\kappa+\tau+1$.
II-D Dedicated Quantum Attack on Constructions Based on XOR-Type Function
The XOR-type function is a class of hidden periodic functions with a specific algebraic structure. For a function $f$, there exists a unique key $k_{0}$ such that $f(k_{0},x)=f(k_{0},x\oplus s)$. Moreover, $f(k,x)$ can be decomposed into two components: $g_{1}(x)$, independent of the key $k$, and $g_{2}(x\oplus\alpha_{k})$, dependent on the input $x$ and the secret state $\alpha_{k}$. The XOR-type function is expressed as
$f(k,x)=g_{1}(x)\oplus g_{2}(x\oplus\alpha_{k}),$
where $\alpha_{k}$ is an $n$-bit secret state derived from the key $k$.
A new function $F^{\mathcal{L}}$ is constructed by summing the value of $f$ over the $t$-dimensional linear subspace $\mathcal{L}=\{u:u=0^{n-t}\|*\}$. The function $F^{\mathcal{L}}$ is written as
$F^{\mathcal{L}}(k,x^{l})=\bigoplus_{u\in\mathcal{L}}f(k,(x^{l}\|0^{t})\oplus u).$
Using the truncated technique, the $n$-bit parameters are decomposed into (nt)(n-t)-bit parts and $t$-bit parts: $X=X^{l}\|X^{r}$, where $X\in\{\alpha_{k},x,s\}$, $|X^{l}|=n-t$, and $|X^{r}|=t$. The function $F^{\mathcal{L}}$ exhibits an (nt)(n-t)-bit truncated period $s^{l}$ when $\alpha_{k}^{l}=\alpha_{k_{0}}^{l}$, as detailed in Property 1 [shi2024dedicated].
Property 1
(Adapted from [shi2024dedicated]) Let $\mathcal{L}=\{u:u=0^{n-t}\|*\}$ and $f(k,x)$ be an XOR-type function such that $f(k,\cdot)$ has a non-trivial period $s$ when $k=k_{0}$. Then $F^{\mathcal{L}}(k,x^{l})=\bigoplus_{u\in\mathcal{L}}f(k,x^{l}\|0^{t}\oplus u)$ has an (nt)(n-t)-bit period $s^{l}$ for $\alpha^{l}_{k}=\alpha^{l}_{k_{0}}$.
In particular, if $f(k,x)=g_{1}(x)\oplus g_{2}(k\oplus x)$, then $F^{\mathcal{L}}(k^{l},x^{l})=\bigoplus_{u\in\mathcal{L}}f(k^{l},(x^{l}\|0^{t})\oplus u)$ has an (nt)(n-t)-bit period $s^{l}$ for $k^{l}=k_{0}^{l}$.
The dedicated quantum attack on constructions based on XOR-type functions using truncated technique was proposed by Shi et al. [shi2024dedicated], and is summarized in Algorithm 3 and Theorem 3 based on the original description.
Algorithm 3 Dedicated Attack on Constructions Based on XOR-Type Function Using Truncated Technique [shi2024dedicated]
Require: $\kappa,n,m,c^{\prime},t,r$. Let $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$ be an XOR-type function to which we have quantum oracle access;
Ensure: the high (nt)(n-t) bits of $k_{0}$ and $s$, namely $k_{0}^{l}$ and $s^{l}$;
Prepare $(\kappa-t)+(n-t)c^{\prime}+mc^{\prime}$ qubits registers $|0^{\kappa-t}\rangle|0^{(n-t)c^{\prime}}\rangle|0^{mc^{\prime}}\rangle$.
1:Apply $(H^{\otimes\kappa-t}\otimes H^{\otimes(n-t)c^{\prime}}\otimes I_{2^{mc^{\prime}}})$ to obtain
[…]
3:Apply $U_{F\mathcal{L}}$ to get
[…]
5:Apply $I_{2^{\kappa-t}}\otimes H^{\otimes(n-t)c^{\prime}}\otimes I_{2^{mc^{\prime}}}$ to compute
6:
$|\psi\rangle=\sum\limits_{\begin{subarray}{c}k^{l}\in\{0,1\}^{\kappa-t},\\ x_{1}^{l},...,x_{c^{\prime}}^{l}\in\{0,1\}^{n-t},\\ v_{1},...,v_{c^{\prime}}\in\{0,1\}^{n-t}\end{subarray}}|k^{l}\rangle$ $(-1)^{x_{1}^{l}\cdot v_{1}}|v_{1}\rangle\ldots(-1)^{x_{c^{\prime}}^{l}\cdot v_{c^{\prime}}}|v_{c^{\prime}}\rangle$
$|F^{\mathcal{L}}(k^{l},x_{1}^{l})\rangle,\ldots,|F^{\mathcal{L}}(k^{l},x_{c^{\prime}}^{l})\rangle.$
7:Repeat Grover iteration $D_{|\psi\rangle}O_{test}$ for $r$ times to get $|\phi\rangle$:
[…]
9:Measure the first $((\kappa-t)+(n-t)c^{\prime})$-bit of $|\phi\rangle$ to get $k_{0}^{l}$ and $v_{1},\cdots,v_{c^{\prime}}$;
10:Compute the period $s^{l}$ by solving the equation system $s^{l}\cdot v_{i}=0$, $i=1,2,\ldots,c^{\prime}$;
Theorem 3
(Adapted from [shi2024dedicated]) Let $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$ be an XOR-type function, there exists a quantum attack that uses $2^{(\kappa-t)/2}$ iterations with $c^{\prime}$ quantum queries per iteration. The probability of falling below a success rate of
$1-2^{-\tau}-\left(2^{-\tau/2-1}+2^{-\tau}+2^{-(\kappa-t)/2+1}\right)^{2}$ (2)
is lower than $2^{n+\kappa-2t-\frac{2^{n-t}}{4c^{\prime}}}$, where $0<t<n$, $c^{\prime}=n+\kappa-2t+\tau+1$.
The quantum complexity of this attack is analyzed as follows. A larger truncation parameter $t$ results in fewer iterations of $2^{(\kappa-t)/2}$ but yields a shorter recovered key of (nt)(n-t)-bit, which consists of $\alpha_{k_{0}}^{l}$ and the period $s^{l}$. Each iteration requires $c^{\prime}$ queries to the function $F^{\mathcal{L}}$. Since $F^{\mathcal{L}}$ is a sum over $2^{t}$ parallel evaluations of $f$, one query to $F^{\mathcal{L}}$ is counted as $2^{t}$ quantum queries to the underlying function $f$. Consequently, the total quantum query complexity to the encryption oracle is $O(2^{(\kappa-t)/2}\cdot 2^{t}\cdot(n-t))=O(2^{(\kappa+t)/2}\cdot(n-t))$.
The remaining $t$-bit $k^{r}$ is then recovered using the Grover-meets-Simon algorithm on the function $f^{\prime}(k^{r},x)$, defined as $f^{\prime}(k^{r},x)=f(k^{l}_{0}\|k^{r},x)$. This function exhibits a period $s$ when the key $k^{l}_{0}\|k^{r}$ is correct. Crucially, this part requires only $O(2^{t/2})$ iterations.
III Offline Dedicated Quantum Attack on Block Ciphers by Constructing Decoupled XOR-Type Function
In this section, we propose an offline dedicated quantum attack on block ciphers by constructing decoupled XOR-type functions. We begin by presenting a new instance based on an XOR-type function for an online attack, and then we illustrate the obstacle encountered when converting it into an offline attack. Subsequently, we propose an offline attack on other instantiations of TPP-PRFs, which can be constructed as decoupled XOR-type functions.
III-A A New Instance for Online Attack and Obstacle to Offline Conversion
In this subsection, we identify PolyMAC, a BBB MAC that is subject to the dedicated quantum attack in Ref. [shi2024dedicated]. Subsequently, we analyze the obstacle that prevents converting this attack to an offline setting.
III-A1 Quantum Attack on PolyMAC with Online Queries
PolyMAC [kim2020tight]. PolyMAC scheme is a Double-block Hash-then-Sum construction based on polynomial evaluation, proposed in 2020. It uses two hashing keys $k_{1},k_{3}\in\{0,1\}^{n}$ and two encryption keys $k_{2},k_{4}\in\{0,1\}^{m}$. We consider the case with the two-block message as follows:
$\text{PolyMAC}(m_{1},m_{2})=E_{k_{2}}(k_{1}^{2}m_{1}\oplus k_{1}m_{2})\oplus E_{k_{4}}(k_{3}^{2}m_{1}\oplus k_{3}m_{2})$ (3)
where $M=m_{1}\|m_{2}$, and $|m_{1}|=|m_{2}|=n$.
Next, we describe the construction of an XOR-type function based on PolyMAC. Let $\beta_{0}$, $\beta_{1}$ be two fixed strings in $\{0,1\}^{n}$, $\beta_{0}\neq\beta_{1}$, and $\alpha_{k}\in\{0,1\}^{n}$, we define the function as
$f:\{0,1\}^{n}\times\{0,1\}^{n}$ $\rightarrow\{0,1\}^{n}$ (4)
$(\alpha_{i},x)$ $\mapsto\text{PolyMAC}(\beta_{0},x)$
$\oplus\text{PolyMAC}\left(\beta_{1},x\oplus\alpha_{i}\right).$
The period $s$ is $(k_{1}\oplus k_{3})(\beta_{0}\oplus\beta_{1})$ when $\alpha_{i}=k_{1}(\beta_{0}\oplus\beta_{1})$ or $\alpha_{i}=k_{3}(\beta_{0}\oplus\beta_{1})$. From Eq. (4), the secret state $\alpha_{i}$ is XORed with $x$ in the second term $\text{PolyMAC}(\beta_{1},x\oplus\alpha_{i})$ of $f(\alpha_{i},x)$.
Now, we propose a new function $F^{\mathcal{L}}$ using truncated input. Let $\mathcal{L}=\{u:u=0^{n-t}\|*\}$, $\alpha_{i}=\alpha_{i}^{l}\|\alpha_{i}^{r}$, $x=x^{l}\|x^{r}$, $|\alpha_{i}^{l}|=|x^{l}|=n-t$, and $|\alpha_{i}^{r}|=|x^{r}|=t$.
$F^{\mathcal{L}}:\{0,1\}^{n-t}\times\{0,1\}^{n-t}$ $\rightarrow\{0,1\}^{m}$ (5)
$(\alpha_{i}^{l},x^{l})$ $\mapsto\bigoplus_{u\in\mathcal{L}}f(\alpha_{i}^{l},x^{l}\|0^{t}\oplus u).$
The function $F^{\mathcal{L}}$ has a period $s^{l}=(k_{1}\oplus k_{3})^{l}(\beta_{0}\oplus\beta_{1})^{l}$ when $\alpha_{i}^{l}=k_{1}^{l}(\beta_{0}\oplus\beta_{1})^{l}$ or $k_{3}^{l}(\beta_{0}\oplus\beta_{1})^{l}$. By applying Algorithm 3, we can recover the high (nt)(n-t)-bits $s^{l}$ and $\alpha_{i}^{l}$.
Taking $\alpha_{i}^{l}=k_{1}^{l}(\beta_{0}\oplus\beta_{1})^{l}$ as an example, we define $f^{\prime}(\alpha_{i}^{r},x)=f(k_{1}^{l}(\beta_{0}\oplus\beta_{1})^{l}\|\alpha_{i}^{r},x)$. When $\alpha_{i}^{r}=k_{1}^{r}(\beta_{0}\oplus\beta_{1})^{r}$, $f^{\prime}(\alpha_{i}^{r},x)=f^{\prime}(\alpha_{i}^{r},x\oplus s)$, which requires $O(2^{t/2})$ quantum queries to $f$.
III-A2 Obstacle to Offline Conversion
Although Section III-A1 showed that PolyMAC is vulnerable to the attack proposed in Ref. [shi2024dedicated], this attack is limited by its requirement for quantum queries to the encryption oracle.
The obstacle to converting this dedicated quantum attack from an online to an offline setting is the tightly coupled property of an XOR-type function $f(i,x)=g_{1}(x)\oplus g_{2}(x\oplus\alpha_{i})$. The issue is that both $g_{1}$ and $g_{2}$ require queries to the encryption oracle. This construction of $f(i,x)$ binds the entire query process to interactive online queries, thus precluding the possibility of offline computation.
Specifically, consider the construction of an XOR-type function based on PolyMAC given by Eq. (4), both the first term $g_{1}(x)=\text{PolyMAC}(\beta_{0},x)$ and the second term $g_{2}(x\oplus\alpha_{i})=\text{PolyMAC}(\beta_{1},x\oplus\alpha_{i})$ of $f(\alpha_{i},x)$ require quantum queries to the encryption oracle. We refer to this as a coupled XOR-type function. Similarly, all XOR-type functions constructed by several BBB MACs, which are proposed by Shi et al. [shi2024dedicated], are coupled XOR-type functions.
Therefore, we introduce a decoupled XOR-type function to enable an offline attack.
Decoupled XOR-type function. Let $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\rightarrow\{0,1\}^{m}$ be an XOR-type function of the form $f(i,x)=g_{1}(x)\oplus g_{2}(x\oplus\alpha_{i})$, where $\alpha_{i}$ is the $n$-bit secret state determined by $i$. If the second term $g_{2}(x\oplus\alpha_{i})$ of $f(i,x)$ is a public function that enables offline computation, we call $f(i,x)$ a decoupled XOR-type function. To highlight its decoupled property, we denote it as an XOR-type function with a public function ($p$-XOR-type function) and express it as $f(i,x)=g_{1}(x)\oplus p(x\oplus\alpha_{i})$, where $p(x\oplus\alpha_{i})$ is a public function.
To facilitate discussion, we introduce the following notation for the case where $i$ is XORed on $x$. Let $\mathcal{L}=\{u:u=0^{n-t}||*\}$. We define a new periodic function $F^{\mathcal{L}}(i^{l},x^{l})$ as follows: $F^{\mathcal{L}}(i^{l},x^{l})=\bigoplus\limits_{u\in\mathcal{L}}f(i^{l},x^{l}||0^{t}\oplus u)=G_{1}^{\mathcal{L}}(x^{l})\oplus P^{\mathcal{L}}(i^{l},x^{l})$, where $G_{1}^{\mathcal{L}}(x^{l})=\bigoplus\limits_{u\in\mathcal{L}}g_{1}(x^{l}\|0^{t}\oplus u)$. Here, $P^{\mathcal{L}}$ is defined based on the public function $p(i,x)$ as $P^{\mathcal{L}}(i^{l},x^{l})=\bigoplus\limits_{u\in\mathcal{L}}p((x^{l}\oplus i^{l})\|0^{t}\oplus u)$. The $p$-XOR-type function is a special class of the XOR-type function. Consequently, Property 1 and Algorithm 3 proposed by Shi et al. [shi2024dedicated] are also applicable to the $p$-XOR-type function.
III-B Other Instances for Offline Attack
Given that the decoupled property of XOR-type functions is crucial for implementing an offline attack, we next explore TPP-PRFs, whose periodic functions naturally possess this property.
III-B1 Instantiation of TPP-PRFs
We show how to construct $p$-XOR-type functions from instantiations of TPP-PRFs.
TPP-PRFs [chen2019build] TPP-PRFs, which use $n$-bit input $x$ and are constructed with two parallel permutations $P_{1}$ and $P_{2}$, achieve $2n/3$ bits of security in the classical setting, defined as follows.
$g:\{0,1\}^{n}$ $\rightarrow\{0,1\}^{n}$ (6)
$x$ $\mapsto l_{33}P_{1}(l_{13}(x)\oplus l_{14}(k_{1}))$
$~~~~~\oplus l_{34}P_{2}(l_{23}(x)\oplus l_{24}(k_{2}))\oplus e(x)\oplus C$
where $e(x)=l_{31}l_{11}(x)\oplus l_{32}l_{21}(x)$, the constant term is $C=l_{31}l_{12}(k_{1})\oplus l_{32}l_{22}(k_{2})\oplus l_{35}(k_{3})\oplus l_{36}(k_{4})$, and $l_{ij}\neq 0$ for $i=1,2,3$, $j=3,4$.
Based on TPP-PRFs, we can construct the following function that is a $p$-XOR-type function.
$f:\{0,1\}^{n}\times\{0,1\}^{n}$ $\rightarrow\{0,1\}^{n}$ (7)
$(i,x)$ $\mapsto g_{1}(x)\oplus p(i,x)$
where $g_{1}(x)=g(x)\oplus l_{33}P_{1}\left(l_{13}(x)\right)\oplus e(x)$, and the public function $p:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\{0,1\}^{n}$, that is, $p(i,x)=l_{34}P_{2}\left(l_{23}(x\oplus i)\right)$. It can be verified that $f(i,x)=f(i,x\oplus l_{13}^{-1}l_{14}(k_{1}))$ when $i=l_{23}^{-1}l_{24}(k_{2})$ or $i=l_{23}^{-1}l_{24}(k_{2})\oplus l_{23}l_{13}^{-1}l_{14}(k_{1})$ in Ref. [sun2025quantum]. Since $i$ is XORed with $x$ in the public function $p(i,x)=l_{34}P_{2}(l_{23}(x\oplus i)$ of $f(i,x)$, the function $f(i,x)$ is a $p$-XOR-type function.
For $p$-XOR-type functions constructed by TPP-PRFs, we can apply truncation techniques to obtain a new periodic function $F^{\mathcal{L}}$.
Let $\mathcal{L}=\{u:u=0^{n-t}\|*\}$. We decompose the (nt)(n-t)-bit key $k_{1}^{l}$ into two parts: $k_{1}^{l_{1}}$ with $p$ bits and $k_{1}^{l_{2}}$ with (ntp)(n-t-p)-bits, where $0\leq p\leq n-t$, and $t$ is a tunable positive integer truncation parameter. We define $G_{1}^{\mathcal{L}}:\{0,1\}^{p}\rightarrow\{0,1\}^{n}$ and the public function $P^{\mathcal{L}}:\{0,1\}^{2n-2t-p}\times\{0,1\}^{p}\rightarrow\{0,1\}^{n}$ as
$G_{1}^{\mathcal{L}}(x^{l_{1}})$ $=\bigoplus\limits_{u\in\mathcal{L}}g(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)$
$~~~~\oplus e(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u),$
$P^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})$ $=\bigoplus\limits_{u\in\mathcal{L}}l_{33}P_{1}\left(l_{13}(x^{l_{1}}\|j^{l_{2}})\|0^{t}\oplus u\right)$
$~~~~\oplus l_{34}P_{2}\left(l_{23}\left((x^{l_{1}}\|0^{n-t-p})\oplus i^{l}\right)\|0^{t}\oplus u\right).$
We construct the new periodic function $F^{\mathcal{L}}:\{0,1\}^{n-t}\times\{0,1\}^{n-t-p}\times\{0,1\}^{p}\rightarrow\{0,1\}^{n}$ as follows:
$F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=$ $G_{1}^{\mathcal{L}}(x^{l_{1}})\oplus P^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}}),$ (8)
where $l_{1}\in\{0,1\}^{p}$, $l_{2}\in\{0,1\}^{n-t-p}$, and $l=l_{1}\|l_{2}$.
This function has the period $s^{l_{1}}=l_{13}^{-1}l_{14}k_{1}^{l_{1}}$ when $i^{l}\|j^{l_{2}}=l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$ (or $i^{l}\|j^{l_{2}}=l_{23}^{-1}l_{24}k_{2}^{l}\oplus l_{23}l_{13}^{-1}l_{14}k_{1}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$). Take $i^{l}\|j^{l_{2}}=l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$ as an example
$F^{\mathcal{L}}(l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}},x^{l_{1}})$
$\qquad=$ $\bigoplus\limits_{u\in\mathcal{L}}l_{33}P_{1}\left((l_{13}x^{l_{1}}\oplus l_{14}k_{1}^{l_{1}})\|l_{14}k_{1}^{l_{2}}\|0^{t}\oplus u\right)$
$\oplus l_{34}P_{2}\left((l_{23}x^{l_{1}}\oplus l_{24}k_{2}^{l_{1}})\|l_{24}k_{2}^{l_{2}}\|0^{t}\oplus u\right)\oplus C$
$\oplus l_{33}P_{1}\left(l_{13}(x^{l_{1}}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}})\|0^{t}\oplus u\right)$
$\oplus l_{34}P_{2}(l_{23}(x^{l_{1}}\|0^{n-t-p}\oplus l_{23}^{-1}l_{24}k_{2}^{l})\|0^{t}\oplus u)$
$\qquad=$ $\bigoplus\limits_{u\in\mathcal{L}}l_{33}P_{1}(l_{13}x^{l_{1}}\|l_{14}k_{1}^{l_{2}}\|0^{t}\oplus u)$
$\oplus l_{34}P_{2}\left((l_{23}x^{l_{1}}\oplus l_{23}l_{13}^{-1}l_{14}k_{1}^{l_{1}}\oplus l_{24}k_{2}^{l_{1}})\|l_{24}k_{2}^{l_{2}}\|0^{t}\oplus u\right)$
$\oplus l_{33}P_{1}\left((l_{13}(x^{l_{1}}\oplus l_{13}^{-1}l_{14}k_{1}^{l_{1}})\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}))\|0^{t}\oplus u\right)\oplus C$
$\oplus l_{34}P_{2}(l_{23}(x^{l_{1}}\oplus l_{13}^{-1}l_{14}k_{1}^{l_{1}})\|0^{n-t-p}\oplus l_{23}^{-1}l_{24}k_{2}^{l})\|0^{t}\oplus u)$
$\qquad=$ $F^{\mathcal{L}}(l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}},x^{l_{1}}\oplus l_{13}^{-1}l_{14}k_{1}^{l_{1}}).$
Next, we demonstrate that several instantiations of TPP-PRFs can be used to construct $p$-XOR-type functions, including $\mathsf{XopEM}$, $\mathsf{SoEM22}$, $\mathsf{SUMPIP}$, and $\mathsf{DS\text{-}SoEM}$.
The Xop construction [bellare1998luby] is defined by the bitwise XOR of the outputs from two distinct pseudorandom permutations (PRPs) applied to the same input $x$:
$\text{Xop}_{E_{1},E_{2}}(x)=E_{1}(x)\oplus E_{2}(x)$ (9)
where $E_{1}$ and $E_{2}$ represent the encryption algorithms of the respective PRPs.
Realizing the Xop construction with two Even-Mansour ciphers, $EM_{1}(x)=P_{1}(x\oplus k_{1})\oplus k_{2}$ and $EM_{2}(x)=P_{2}(x\oplus k_{3})\oplus k_{4}$, resulting in the $\mathsf{XopEM}$ function.
$\mathsf{XopEM}(x)=P_{1}(x\oplus k_{1})\oplus P_{2}(x\oplus k_{3})\oplus k_{2}\oplus k_{4}$ (10)
We can obtain a $p$-XOR-type function as follows: $f(i,x)=P_{1}(x\oplus k_{1})\oplus P_{2}(x\oplus k_{3})\oplus P_{1}(x)\oplus P_{2}(x\oplus i)\oplus k_{2}\oplus k_{4}$, where $g_{1}(x)=\mathsf{XopEM}(x)$, and the public function is $p(i,x)=P_{1}(x)\oplus P_{2}(x\oplus i)$. This function has a period $s=k_{1}$ precisely when $i=k_{3}$.
Similarly, we define $G^{\mathcal{L}}:\{0,1\}^{p}\rightarrow\{0,1\}^{n}$ and $P^{\mathcal{L}}:\{0,1\}^{n-t}\times\{0,1\}^{n-t-p}\times\{0,1\}^{p}\rightarrow\{0,1\}^{n}$, where $l_{1}\in\{0,1\}^{p}$, $l_{2}\in\{0,1\}^{n-t-p}$, and $l=l_{1}\|l_{2}$:
$G^{\mathcal{L}}(x^{l_{1}})=$ $\bigoplus\limits_{u\in\mathcal{L}}\mathsf{XopEM}(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u),$
$P^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=$ $\bigoplus\limits_{u\in\mathcal{L}}P_{1}(x^{l_{1}}\|j^{l_{2}}\|0^{t}\oplus u)$
$\oplus P_{2}(((x^{l_{1}}\|0^{n-t-p})\oplus i^{l})\|0^{t}\oplus u).$
The periodic function $F^{\mathcal{L}}:\{0,1\}^{n-t}\times\{0,1\}^{n-t-p}\times\{0,1\}^{p}\rightarrow\{0,1\}^{n}$ is defined as:
$F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=G^{\mathcal{L}}(x^{l_{1}})\oplus P^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}}).$ (11)
It can be verified that $F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}}\oplus k_{1}^{l_{1}})$ when $i^{l}\|j^{l_{2}}=k_{3}^{l}\|k_{1}^{l_{2}}$ or $i^{l}\|j^{l_{2}}=k_{1}^{l}\oplus k_{3}^{l}\|k_{1}^{l_{2}}$.
SoEM22 [chen2019build]. The Sum of Even-Mansour (SoEM) construction employs two public permutations $P_{1}$, $P_{2}$, and two $n$-bit keys $k_{1}$, $k_{2}$. It is a specific instance of the XOR construction where $k_{3}=k_{2}$ and $k_{4}=k_{1}$, defined as:
$\mathsf{SoEM22}(x)=P_{1}(x\oplus k_{1})\oplus P_{2}(x\oplus k_{2})\oplus k_{1}\oplus k_{2}.$ (12)
We can construct a $p$-XOR-type function $f(i,x)=\mathsf{SoEM22}(x)\oplus P_{1}(x)\oplus P_{2}(x\oplus i)$, where $P_{1}(x)\oplus P_{2}(x\oplus i)$ is a public function. This function satisfies $f(i,x)=f(i,x\oplus k_{1})$ when $i=k_{2}$. The new periodic function $F^{\mathcal{L}}$ is defined as:
$F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=$ $\bigoplus\limits_{u\in\mathcal{L}}\mathsf{SoEM22}(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)$ (13)
$\oplus P_{1}(x^{l_{1}}\|j^{l_{2}}\|0^{t}\oplus u)$
$\oplus P_{2}(((x^{l_{1}}\|0^{n-t-p})\oplus i^{l})\|0^{t}\oplus u).$
This function has a period $k_{1}^{l_{1}}$ when $i^{l}\|j^{l_{2}}=k_{2}^{l}\|k_{1}^{l_{2}}$ or $i^{l}\|j^{l_{2}}=k_{1}^{l}\oplus k_{2}^{l}\|k_{1}^{l_{2}}$.
If $P_{1}=P$ and $P_{2}=P^{-1}$, the above $\mathsf{SoEM22}$ construction corresponds to $\mathsf{SUMPIP}$ [guo2019beyond]. In the same way, a $p$-XOR-type function $f(i,x)$ and a new periodic function $F^{\mathcal{L}}$ can be constructed based on the truncation parameter.
DS-SoEM [bhattacharjee2022cencpp]. This construction is a domain-separated variant of the Xop construction that uses a $d$-bit domain-separation constant and is a sum of two Even-Mansour ciphers. For an input $x\in\{0,1\}^{n-d}$, it is defined as:
$\mathsf{DS\text{-}SoEM}(x)=$ $P\left(\left(x\oplus\mathrm{msb}_{n-d}(k_{1})\right)\|0^{d}\right)$ (14)
$\oplus P\left(\left(x\oplus\mathrm{msb}_{n-d}(k_{2})\right)\|1^{d}\right)\oplus k_{1}\oplus k_{2},$
where $\mathrm{msb}_{n-d}$’ denotes truncation to the ndn-d most significant bits.
We define a $p$-XOR-type function $f(i,x)=\mathsf{DS\text{-}SoEM}(x)\oplus P(x\|0^{d})\oplus P\left((x\oplus i)\|1^{d}\right)$, where the public function is $P(x\|0^{d})\oplus P\left((x\oplus i)\|1^{d}\right)$. This function satisfies $f(i,x)=f\left(i,x\oplus\mathrm{msb}_{n-d}(k_{1})\right)$ when $i=\mathrm{msb}_{n-d}(k_{2})$. The new function $F^{\mathcal{L}}$ is defined as:
$F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=$ $\bigoplus_{u\in\mathcal{L}}\mathsf{DS\text{-}SoEM}(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)$ (15)
$\oplus P\left((x^{l_{1}}\|j^{l_{2}}\|0^{t}\oplus u)\|0^{d}\right)$
$\oplus P\left(\left((x^{l_{1}}\|0^{n-t-p}\oplus i^{l})\|0^{t}\oplus u\right)\|1^{d}\right).$
This function is periodic, with $F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}}\oplus[\mathrm{msb}_{n-d}(k_{1})]^{l_{1}})$ when $i^{l}\|j^{l_{2}}=[\mathrm{msb}_{n-d}(k_{2})]^{l}\|[\mathrm{msb}_{n-d}(k_{1})]^{l_{2}}$.
Algorithm 4 Offline Dedicated Attack on block cipher constructions based on $p$-XOR-Type Function Using Truncated Technique
1:
2:Prepared $\left|\psi_{G_{1}^{\mathcal{L}}}\right\rangle=\otimes^{c^{\prime}}\left(\sum_{x^{l}\in\{0,1\}^{n-t}}|x^{l}\rangle|G_{1}^{\mathcal{L}}(x^{l})\rangle\right)$,
3:Let $\mathcal{L}=\{u:u=0^{n-t}||*\}$, $f(i,x)$ be a $p$-XOR-type function, and $f(i,\cdot)$ have a non-trivial period $s$ when $i=i_{0}$.
4:
5:The high (kt)(k-t)-bit of $i_{0}$ and (nt)(n-t)-bit of $s$, namely $i_{0}^{l}$ and $s^{l}$, s.t., $F^{\mathcal{L}}(i^{l},x^{l})=F^{\mathcal{L}}(i^{l},x^{l}\oplus s^{l})$ when $i^{l}=i^{l}_{0}$.
6:Start $\left|\psi_{G_{1}^{\mathcal{L}}}\right\rangle|0^{\kappa-t}\rangle$.
7:Apply $I_{c^{\prime}(n-t+m)}\otimes H^{\otimes\kappa-t}$ to obtain $|\Psi\rangle$:
[…]
9:Repeat $\mathcal{O}(2^{(\kappa-t)/2})$ Grover iterations:
[…]
11:Note that the test oracle is a unitary operator that takes $\left|\psi_{G_{1}^{\mathcal{L}}}\right\rangle|i^{l}\rangle$ as input, and tests whether $F^{\mathcal{L}}(i^{l},\cdot)$ has a hidden period (see Algorithm 5 for details).
12:Measure the index $i^{l}$ to obtain $i_{0}^{l}$.
III-B2 Offline Dedicated Quantum Attacks on Block Ciphers Based on TPP-PRFs
Based on the above $p$-XOR-type functions constructed by TPP-PRFs, we propose an offline dedicated quantum attack on these constructions, utilizing the truncated technique in the Q1 (Q2) model to reduce query complexity to the encryption oracle.
We present a new algorithm for an offline dedicated quantum attack on constructions based on TPP-PRFs, which integrates the offline Simon’s algorithm with the dedicated quantum attack on block ciphers by constructing $p$-XOR-type functions. Our algorithm is separated into two phases: the online phase prepares the superposition state $|\psi_{G_{1}^{\mathcal{L}}}\rangle$ via classical or quantum queries to the encryption oracle $U_{g_{1}}$ (requiring $O(2^{n})$ classical queries in the Q1 model by Algorithm 6, or $O(2^{t}\cdot(n-t))$ quantum queries in the Q2 model by Algorithm 7), and the offline computation phase performs the remaining quantum computations independently of the encryption oracle (using the test procedure in Algorithm 5 to check whether the function $F^{\mathcal{L}}$ has a period without any new queries to the encryption oracle). This separation achieves a reduction in query complexity, as Grover iterations are executed offline after a limited number of online interactions.
The offline dedicated quantum attack on constructions based on $p$-XOR-type functions in the Q1 (Q2) model is presented in Algorithm 4.
The overall quantum attack aims to find the high $(\kappa-t)$ bits $i_{0}^{l}$ such that it $F^{\mathcal{L}}(i_{0}^{l},\cdot)$ exhibits a non-trivial period $s^{l}$. Starting with the superposition state $\left|\psi_{G_{1}^{\mathcal{L}}}\right\rangle=\otimes^{c^{\prime}}\left(\sum_{x^{l}\in\{0,1\}^{n-t}}|x^{l}\rangle|G_{1}^{\mathcal{L}}(x^{l})\rangle\right)$, the input is $\left|\psi_{G_{1}^{\mathcal{L}}}\right\rangle|0^{\kappa-t}\rangle$ of $(\kappa-t)+c^{\prime}(n-t)+c^{\prime}m$ qubits. The initial state $|\Psi\rangle$ required for the Grover iteration is obtained by several $H$ transforms. The iterative function $G=D_{|\Psi\rangle}O_{\text{test}}$ is run for $\mathcal{O}(2^{(\kappa-t)/2})$ iterations, where $D_{|\Psi\rangle}=(2|\Psi\rangle\langle\Psi|-I)$. The final step is measuring the index $i^{l}$ to obtain $i_{0}^{l}$. The implementation of $U_{P^{\mathcal{L}}}$ and $O_{\text{test}}$ in Algorithm 4 is similar to that in Ref. [shi2024dedicated]. Specifically, $U_{P^{\mathcal{L}}}$ implements the function
$P^{\mathcal{L}}(i^{l},x^{l})=\bigoplus_{u\in\mathcal{L}}p((x^{l}\oplus i^{l})\|0^{t}\oplus u),$
which is the XOR sum of $2^{t}$ public functions $p$ and can be implemented in parallel without a sequential relationship, where $t>0$ is a tunable truncation parameter.
Algorithm 5 The procedure checks whether a function $F^{\mathcal{L}}(i^{l},x^{l})=G_{1}^{\mathcal{L}}(x^{l})\oplus P^{\mathcal{L}}(i^{l},x^{l})$ has a period, without making any new queries to $g_{1}$.
1:
2:Prepared $|\psi_{G_{1}^{\mathcal{L}}}\rangle$.
3:Start $|\psi_{G_{1}^{\mathcal{L}}}\rangle|b\rangle$:
[…]
5:Apply $c^{\prime}$ $U_{P^{\mathcal{L}}}$ to obtain $|\psi_{F^{\mathcal{L}}}\rangle$:
6:
$|\psi_{F^{\mathcal{L}}}\rangle=$ $\otimes^{c^{\prime}}\left(\sum_{x^{l}\in\{0,1\}^{n-t}}|x^{l}\rangle|G_{1}^{\mathcal{L}}(x^{l})\oplus P^{\mathcal{L}}(x^{l})\rangle\right)|b\rangle$
$=$ $\otimes^{c^{\prime}}\left(\sum_{x^{l}\in\{0,1\}^{n-t}}|x^{l}\rangle|F^{\mathcal{L}}(x^{l})\rangle\right)|b\rangle.$
7:Apply $(H^{\otimes(n-t)}\otimes I_{m})^{\otimes c^{\prime}}\otimes I_{1}$ to $|\psi_{F^{\mathcal{L}}}\rangle$, to get:
8:
$\left(\sum\limits_{v_{1},x_{1}^{l}\in\{0,1\}^{n-t}}(-1)^{v_{1}\cdot x_{1}^{l}}|v_{1}\rangle|F^{\mathcal{L}}(x_{1}^{l})\rangle\right)\otimes\cdots$
$\otimes\left(\sum\limits_{v_{c^{\prime}},x_{c^{\prime}}^{l}\in\{0,1\}^{n-t}}(-1)^{v_{c^{\prime}}\cdot x_{c^{\prime}}^{l}}|v_{c^{\prime}}\rangle|F^{\mathcal{L}}(x_{c^{\prime}}^{l})\rangle\right)|b\rangle.$
9:Compute $d:=\dim(\text{Span}(v_{1},\dots,v_{c^{\prime}}))$.
10:if $d=n$ then
11:  set $r:=0$
12:else
13:  set $r:=1$, and add $r$ to $b$
14:end if
15:Uncompute steps 3-2 to obtain:
[…]
To achieve a usable success probability after $r$ iterations, we need $c^{\prime}=n+\kappa-2t+\pi+1$ queries per iteration. We now give the complexity and success probability of the offline dedicated quantum attack on constructions based on $p$-XOR-type functions in the Q1 (Q2) model in Theorem 4.
Theorem 4
Let $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$ be a decoupled XOR-type function, then there exists an offline dedicated quantum attack on constructions based on TPP-PRFs in Algorithm 4. This attack can recover $i_{0}^{l}\in\{0,1\}^{\kappa-t}$ with a time complexity of $O(2^{(\kappa-t)/2}\cdot n^{3})$, requiring $O(2^{n})$ classical queries to the encryption oracle in the Q1 model, and $O(2^{t}\cdot(n-t))$ quantum queries in the Q2 model. The probability of falling below a success rate of
$1-2^{-\tau}-\left(2^{-\tau/2-1}+2^{-\tau}+2^{-(\kappa-t)/2+1}\right)^{2}$ (16)
is lower than $2^{n+\kappa-2t-\frac{2^{n-t}}{4c^{\prime}}}$, where $0<t<n$, $c^{\prime}=n+\kappa-2t+\tau+1$.
For $p$-XOR-type functions and their new periodic functions $F^{\mathcal{L}}$ constructed from instantiations of TPP-PRFs in Section III-B1, we can implement offline dedicated attacks to recover the period of $F^{\mathcal{L}}$ by Theorem 4.
Algorithm 6 Prepare $|\psi_{G_{1}^{\mathcal{L}}}\rangle$ in the Q1 model
1:
2:Classical query to the encryption oracle $g_{1}$.
[…]
4:Apply $(H^{\otimes(n-t)}\otimes I_{m})^{\otimes c^{\prime}}$ to obtain:
[…]
6:For each $x^{l}\in\{0,1\}^{n-t}$, classical queries to $g_{1}$ to get $\bigoplus\limits_{u\in\mathcal{L}}g_{1}(x^{l}||0^{t}\oplus u)$.
7:Apply unitary operation, which writes $\bigoplus\limits_{u\in\mathcal{L}}g_{1}(x^{l}||0^{t}\oplus u)$ in the second register if the first contains the value $x^{l}$:
8:
$|\psi_{G_{1}^{\mathcal{L}}}\rangle$ $=\otimes^{c^{\prime}}\left(\sum_{x^{l}\in\{0,1\}^{n-t}}|x^{l}\rangle|\bigoplus\limits_{u\in\mathcal{L}}g_{1}(x^{l}||0^{t}\oplus u)\rangle\right)$
$=\otimes^{c^{\prime}}\left(\sum_{x^{l}\in\{0,1\}^{n-t}}|x^{l}\rangle|G_{1}^{\mathcal{L}}(x^{l})\rangle\right).$
Algorithm 7 Prepare $|\psi_{G_{1}^{\mathcal{L}}}\rangle$ in the Q2 model
1:
2:quantum queries to the encryption oracle $g_{1}$.
[…]
4:Apply $(H^{\otimes(n-t)}\otimes I_{m})^{\otimes c^{\prime}}$ to obtain:
[…]
6:The XOR sum of $2^{t}$ distinct $U_{g_{1}}$, which can be implemented by using CNOTs, we can obtain $|\psi_{G_{1}^{\mathcal{L}}}\rangle$:
7:
$|\psi_{G_{1}^{\mathcal{L}}}\rangle$ $=\otimes^{c^{\prime}}\left(\sum_{x^{l}\in\{0,1\}^{n-t}}|x^{l}\rangle|G_{1}^{\mathcal{L}}(x^{l})\rangle\right).$
In the Q1 model, this attack recovers the key $l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$ of block ciphers based on TPP-PRFs, requiring $O(2^{p+t})$ classical queries to the encryption oracle and $O(2^{(2n-2t-p)/2}\cdot n^{3})$ offline quantum computation time, where 0<p<nt0<p<n-t, $0<t<n$, $l_{2}\in\{0,1\}^{n-t-p}$ and $l\in\{0,1\}^{n-t}$. Let $D$ be the number of classical queries and $T$ be the time of offline quantum computations. This attack achieves a tradeoff of $T^{2}D=2^{2n-t}$, which is balanced at $T=D=\tilde{O}(2^{(2n-t)/3})$. A larger truncation parameter $t$ reduces both the time and the classical query complexity but results in a shorter recovered key length.
The remaining keys can be recovered as follows. We can recover the period $s^{l_{1}}=l_{13}^{-1}l_{14}k_{1}^{l_{1}}$ by solving the equation system $s^{l_{1}}\cdot v_{i}^{l_{1}}=0$, requiring $O(n^{3})$ offline computation time. Subsequently, the remaining $t$-bit can be recovered by applying the offline Simon’s algorithm on $f^{\prime}(k^{r},x)=f(k_{2}^{l}\|i^{r},x)$, which has the period $s$ when $i^{r}=k_{0}^{r}$.
In the Q2 model, our offline attack proposed in Theorem 4 involves the construction of $F^{\mathcal{L}}$ that differs from in Section III-B1 by making it unnecessary to partition the (nt)(n-t)-bit subkey $k_{1}^{l}$ into $k_{1}^{l_{1}}$ with $p$-bit and $k_{1}^{l_{2}}$ with (ntp)(n-t-p)-bit.
TABLE I: Previous and new quantum attacks on $\mathsf{SoEM22}$
Algorithm Model Iteration Query Time Reference
Grover-meets-Simon [leander2017grover] Q2 $2^{n/2}$ $O(2^{n/2}\cdot n)$ $O(2^{n/2}\cdot n^{3})$ [shinagawa2022quantum]
Dedicated attack by constructing an XOR-type function [shi2024dedicated] Q2 $2^{(n-t)/2}$ $O(2^{(n+t)/2}\cdot(n-t))$ $O(2^{(n-t)/2}\cdot n^{3})$ Ours
Offline dedicated attack by constructing a $p$-XOR-type function Q2 $2^{(n-t)/2}$ $O(2^{t}\cdot(n-t))$ $O(2^{(n-t)/2}\cdot n^{3})$ Ours
Offline Simon’s algorithm [bonnetain2019quantum] Q1 $2^{2n/3}$ $O(2^{2n/3})$ $O(2^{2n/3}\cdot n^{3})$ [sun2025quantum]
Offline dedicated attack by constructing a $p$-XOR-type function Q1 $2^{(2n-t)/3}$ $O(2^{(2n-t)/3})$ $O(2^{(2n-t)/3}\cdot n^{3})$ Ours
  • 1
    The fourth column “Query” indicates classical queries to the encryption oracle in the Q1 model and quantum queries in the Q2 model.
  • 2
    The fifth column “Time” represents the offline quantum computing time complexity in the Q1 model.
Taking the $\mathsf{SoEM22}$ construction as an example, we construct a $p$-XOR-type function $f(i,x)=\mathsf{SoEM22}\oplus P_{1}(x)\oplus P_{2}(x\oplus i)$ with the public function $P_{2}(x\oplus i)$, then obtain a new periodic function $F^{\mathcal{L}}(i^{l},x^{l})$ as follows:
$F^{\mathcal{L}}(i^{l},x^{l})=$ $\bigoplus\limits_{u\in\mathcal{L}}\mathsf{SoEM22}(x^{l}\|0^{t}\oplus u)\oplus P_{1}(x^{l}\|0^{t}\oplus u)$ (17)
$\oplus P_{2}\left((x^{l}\oplus i^{l})\|0^{t}\oplus u\right).$
This function has an (nt)(n-t)-bit period $s^{l}=k^{l}_{1}$ when $i^{l}=k^{l}_{2}$. We first prepare the state $|\psi_{G_{1}^{\mathcal{L}}}\rangle$ by requiring $O(2^{t}\cdot(n-t))$ quantum queries to the $\mathsf{SoEM22}$ oracle by Algorithm 7. Then, we apply the offline attack to successfully recover the period $k^{l}_{1}$ with $O(2^{(n-t)/2}\cdot n^{3})$ offline quantum computation time by Algorithm 4. Compared to the attack by Shi et al. [shi2024dedicated], our method reduces the quantum query complexity from $O(2^{(n+t)/2}\cdot(n-t))$ to $O(2^{t}\cdot(n-t))$ with the same time complexity, demonstrating the advantage of our offline attack in the Q2 model (See Table I for a comparison with previous attacks and ours on the $\mathsf{SoEM22}$ construction).
IV Conclusion
This work successfully analyzed the quantum security of PolyMAC and block cipher structures based on TPP-PRFs by constructing XOR-type functions, thereby resolving the open question raised by Shi et al. [shi2024dedicated]. We introduce a decoupled XOR-type function, termed $p$-XOR-type function, and propose an offline dedicated quantum attack on constructions based on TPP-PRFs. Our offline attack yields significant improvements. In the Q2 model, the quantum query complexity is reduced from $O(2^{(n+t)/2}\cdot(n-t))$ to $O(2^{t}\cdot(n-t))$ with the same time complexity. In the Q1 model, our attack recovers the key of TPP-PRFs by making $O(2^{p+t})$ classical queries and performing $O(2^{(2n-2t-p)/2}\cdot n^{3})$ offline quantum computation time, which balances at $\tilde{O}(2^{(2n-t)/3})$. Our results provide new insights into the quantum security of cryptographic schemes, including TPP-PRFs and specific instantiations such as $\mathsf{XopEM}$, $\mathsf{SoEM22}$, $\mathsf{SUMPIP}$, and $\mathsf{DS\text{-}SoEM}$.
Our results open several promising avenues for future research: First, exploring the quantum security of more cryptographic structures that can construct decoupled XOR-type functions in the Q1 model. Second, developing more generic and effective offline dedicated quantum attacks by identifying algebraic properties applicable to attacking on block ciphers, including $\mathsf{Feistel}$ [dong2018quantum, ito2019quantum, sun2023quantumAQT], $\mathsf{MISTY}$ [canale2022simon], $\mathsf{Lai\text{-}Massey}$ structure [mao2022quantum], and other permutation-based pseudorandom functions (e.g., $\mathsf{EDM}$ [cogliati2016ewcdm] and $\mathsf{EDMD}$ structure [mennink2017encrypted]).
Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant Nos. 62372048, 62272056, 62371069).